### v1

# Sensing with near-infrared laser trapped fluorescent nanodiamonds

Abstract
Biosensing based on optically trapped fluorescent nanodiamonds is an intriguing research direction potentially allowing to resolve biochemical processes inside living cells. Towards this goal, we investigate infrared near (NIR) laser irradiation at 1064 nm on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers. By conducting comprehensive experiments, we aim to understand how NIR exposure influences the fluorescence and sensing properties of FNDs and to determine the potential implications for the use of FNDs in various sensing applications. The experimental setup involved exposing FNDs to varying intensities of NIR laser light and analyzing the resultant changes in their optical and physical properties. Key measurements included T1 relaxation times, optical spectroscopy, and optically detected magnetic resonance (ODMR) spectra. The findings reveal how increased NIR laser power correlates with alterations in ODMR central frequency but also that charge state dynamics under NIR irradiation of NV centers plays a role. We suggest protocols with NIR and green light that mitigate the effect of NIR, and demonstrate that FND biosensing works well with such a protocol.
Nanodiamond, Optical trapping, Temperature sensing, NV-centers
preprint: APS/123-QED
Figure 1: Graphical abstract illustrating the exploration of effects of NIR laser on NV center for bio-sensing.
I Introduction
Fluorescent nanodiamonds (FNDs) containing Nitrogen-Vacancy (NV) centers have attracted significant interest due to their combined optical and magnetic properties, which render them highly versatile for applications in particular in biomedical sensing and imaging [1, 2]. The capacity of NV centers for optical polarization and readout of spin states earmarks them as formidable tools for advanced sensing applications [3]. Yet, the utility of these sensors is heavily influenced by the photodynamics of the NV centers’ spin- and charge-states, which in turn governs their sensitivity and accuracy [4].
NV centers possess an electronic spin S=1 system, characterized by spin sublevels $m_{s}=0$ and $m_{s}=\pm 1$, which are separated by a zero-field splitting due to spin-spin interaction. This zero-field splitting is sensitive to external parameters such as temperature, strain, and electric fields, while the energies of the spin sublevels can be further shifted by magnetic fields through the Zeeman effect [5, 6]. Additionally, the optical spectrum of NV centers, including their fluorescence intensity and wavelength, is determined by the charge state distribution and depends on the local electrochemical potential and intrinsic charge state dynamics. These dependencies allow NV centers to act as versatile nanoscale sensors, providing insights into local magnetic fields across various frequency ranges, temperature fluctuations, and surface electrochemical conditions. Therefore, to accurately measure external parameters like magnetic fields or temperature with high sensitivity in biological environments, it is essential to understand how external control—such as a near infrared trapping laser—affect the intrinsic properties of NV centers [7, 8].
Trapping of FNDs with optical tweezers is an attractive approach as it potentially allows for spatially resolved sensing in challenging environments, with a resolution determined by the size of the FND and of the trapping potential [9]. To achieve this in biological samples and minimize cell and tissue damage, the wavelength of the intense trapping laser must be in the biological transparency window, such as 1064 nm, where residual absorption of tissue components is minimal [10].
In this work, we employed three principal methods to investigate the properties of NV centers in FNDs under NIR light exposure: spin polarization relaxometry (T1-relaxometry) [11, 12], optically detected magnetic resonance (ODMR) spectroscopy [13, 14], and photoluminescence (PL) spectrum analysis [15]. PL spectroscopy involves measuring the emitted light from the NV centers when excited by a laser, providing information on the optical properties and charge states of the NV centers [16]. ODMR combines optical and microwave techniques to detect changes in the NV centers’ fluorescence in response to varying magnetic fields and relative temperature changes, allowing for precise measurement of the NV centers’ electron spin states and local magnetic environment [17]. T1-relaxometry, on the other hand, measures the longitudinal relaxation time (T1), which is the time it takes for the NV centers’ spin states to return to thermal equilibrium after being polarized [18]. This method is valuable for sensing processes and can detect various types of magnetic noise at the spin-transition frequency, such as free radicals [12]. Combined, these techniques provide a comprehensive analysis of the NV centers’ optical, magnetic, and spin properties under varying experimental conditions [4]. We employ a controlled 30 to 60mW NIR laser, with a pulsing sequence inspired by [19], and suggest that the approach can facilitate more precise and reliable intracellular biosensing assessments while keeping any undesirable NIR effect low.
Previous efforts by Aslam et al. [20] have already explored the impact of near-infrared (NIR) laser irradiation on the properties of NV centers, revealing that NIR irradiation can significantly modify NV photoluminescence and enhance spin-lattice relaxation processes, using factors such as NIR laser power (around 30mW) and pulse duration (around 30ns) [21]. The existing literature provides different and sometimes conflicting explanations on the physical mechanisms induced by NIR radiation, which modify both the optical and spin properties of NV centers [22, 15, 9, 21, 23, 14].
In our experiment, we first investigated the effect of the NIR laser on the fluorescence of FNDs in two scenarios: dry FNDs and optically trapped FNDs. We then demonstrated the sensing capabilities of the optically trapped FNDs, which included measurements in different pH environments, controlled temperature variations, and the addition of a paramagnetic species (Gd3+). We propose that changes in spin and photoluminescence properties may be due to three effects: i) heating of the FNDs induced directly by NIR laser irradiation of the FNDs by the NIR laser, ii) a charge state recombination and ionization rate that depends on the NIR laser power and iii) a sensitivity effect, where the NIR laser alters the FNDs’ response to their environment [24]. The sensitivity effect is particularly important for those sensing protocols that rely on NV centers positioned close to the surface of the FNDs [4, 17]. For example, the pH of the surrounding environment plays a significant role in the behavior of NV centers in FNDs. Variations in pH can influence the charge state dynamics of NV centers, which in turn affects their fluorescence properties and overall sensitivity in sensing applications [24, 18, 13, 25].
In this article, we build upon existing hypotheses that the effects of NIR irradiation on NV centers involve both heating, as suggested by previous studies [22], and redistribution between charge states [21, 23]. We aim to provide a more comprehensive understanding of these combined effects, thus offering new insights into optimizing optically trapped NV-based sensing protocols [9]. The paper is organized as follows: Section II describes the methods, including sample preparation, the optical setup, FND trapping, and sensing experiments. Section III presents the results, focusing on photoluminescence spectrum analysis, optically detected magnetic resonance measurements, T1 relaxometry, and a combined statistical analysis. Section IV discusses the implications of the findings, including thermal effects, charge state dynamics, and their interplay, along with alternative considerations. Finally, Section V explores the sensing capabilities of trapped FNDs.
II Methods
II.1 Sample Preparation
Fluorescent FNDs with a nominal diameter of 120 nm (Sigma-Aldrich, 798088) were used for most of the experiments reported in this article if not stated otherwise. They possess a significant number of nitrogen-vacancy (NV) centers due to high concentration of 3ppm. To prepare the FND suspension, 1 µL of the commercial stock solution of FNDs was diluted in 300-700 µL of deionized water depending on the desired concentration and before sonicating for 10 minutes. Next, the FND suspension was deposited on a petri dish (Mattek, part no P35G-1.5-10-C.), thoroughly cleaned to remove any contaminants that could interfere with the experiments and plasma ionised in order to make the surface hydrophilic, ensuring an even distribution of FNDs on the surface. Plasma treatment was applied for 30 seconds in a simple custom-device with a closed chamber in PMMA and under ambient conditions in which an electrode (Electro Technic Product, BD-20V) could be positioned close to the petri dish surface. The plasma treatment modifies the surface energy of the petri dish by breaking chemical bonds on the surface and introducing polar functional groups, thereby inducing the hydrophilic state [26]. Approximately 3 µL of the FND suspension was then carefully deposited onto the clean petri dish using a micropipette. The suspension was gently spread across the surface before allowing it to dry. The drying process took place at room temperature to minimize any potential alterations to the FND properties potentially caused by excessive heating or cooling. After the solution had completely evaporated, the sample was ready for the optical experiments described in the following. The preparation of FNDs in the liquid environment also starts with the dilution steps and cleaning of the petri dish, but to maintain the FNDs in solution, the petri dish was not plasma treated and the solution was not left for evaporation.
FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments [27]. The FNDs (Sigma-Aldrich 1003626896) were prepared following the same protocol. A suspension of FNDs was created by diluting 1 µL of the commercial stock solution in 1000 µL of deionized water, followed by sonication for 10 minutes to ensure a homogeneous dispersion. Approximately 500 µL of the suspension was deposited in a Petri dish that was thoroughly cleaned.
II.2 Optical Setup
The dried FND sample was placed inside a mini-incubator (pecon XS 2000) in an optical setup based on a customized Thorlabs OTKB/M. Customization included a different objective (Nikon, CFI PLAN APO LAMBDA 100x OIL NA 1.45 WD 0.13) and long-distance condenser lens (Mitutoyo M Plan Apo 10x / 0.26). The setup is constructed around an inverted microscope configuration, equipped with a high-resolution camera (Teledyne Prime 95B), and two lasers for the simultaneous illumination and manipulation of the FNDs. A green laser (Coherent Verdi G5) with a wavelength of 532 nm was utilized for the excitation of the NV centers, enabling the initialization and readout of their spin states. The second laser (Cobolt Rumba, 3W) is operating in the near-infrared (NIR) at a wavelength of 1064 nm and used for FND trapping. This wavelength falls within the biological transparency window, which minimizes absorption and reduces the risk of significant heating effects in biological samples. The two lasers were aligned and their beams combined using a dichroic mirror, ensuring they shared the same optical path before being focused onto the sample through the microscope objective, as detailed in Fig. 2. By trapping an FND and then adjusting the position of the green laser to maximize the signal on the photon counter, we ensured the same focal point for both the green and NIR lasers. Additionally, acousto-optic modulators (AOMs) were used to precisely control the timing and intensity of the laser beams, enabling fine-tuned manipulation of the FNDs.
As illustrated in Fig. 2, several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection. A motorized flip-mirror ensures the ability to rapidly shift between camera detection or either spectrometer or photon count detection. The fluorescence spectrum of the FNDs is recorded and analyzed by a fiber coupled spectrometer (Ocean Optics X00428) while alternatively photon counts are measured with a fiber coupled avalanche photo detector (Excelitas SPCM-AQRH-13-Fc). In addition, to measure the ODMR spectrum of FNDs, an Agilent N5181A signal generator connected to an amplifier (MiniCircuits ZHL-45) is connected to a copper wire antenna placed approximately 1mm above the sample. When the flip-mirror is positioned towards the camera, the latter captures high-resolution images of the FNDs illuminated during the experiments in real time. The combination of different detection methods thus made it possible to continuously monitor the position of the FNDs, determine whether they were isolated or in aggregates, and analyse their photoluminescence properties throughout the experimental process.
Figure 2: Illustration of the optical setup with the most important lasers, optical lenses, filters and equipment.
II.3 FND Trapping
For the FND trapping experiments, solution samples were used to ensure the FNDs remained mobile during the experiment.
The NIR laser was utilized for trapping the FNDs in the focused beam [28, 29, 30, 31]. For $T_{1}$ relaxometry, the temporal sequence of the green and trapping lasers, as illustrated in Fig. 3, is carefully designed to minimize the overlap between the two lasers [19]. This approach reduces potential spin-mixing effects from the NIR laser, while maintaining the stability of the trap. The green laser was pulsed for 35 µs for the excitation and 15 µs for the readout of the NV center spin states. The NIR laser was pulsed for 20 µs for the trapping of the FNDs. By using this sequence, we optimized the trapping stability during $T_{1}$ relaxometry measurement and minimized NIR irradiation potential impact on the measurements.
The sensing experiments utilized the FND trapping protocol described below, and the CW NIR laser power was set to 60 mW to stabilize the FNDs in the optical trap. However, for manipulations involving changes to the surrounding medium, the CW NIR laser power was increased to 100 mW to ensure that the FND remained trapped despite turbulence in the medium. After the modifications, the laser power was reduced back to 60 mW for subsequent measurements.
Figure 3: Laser sequence for T1 measurement. In green the 532 nm laser and in red the 1064 nm laser. $t_{NIR}=20\mu s$, $t_{I}=35\mu s$, $t_{RO}=15\mu s$ and $\tau$ from $7.2\mu s$ to $720\mu s$.
II.4 Sensing Experiments
II.4.1 Gd3+ Magnetic Noise Detection
For the experiments detecting paramagnetic species, a gadolinium solution (0.45 M in acetate) was gradually added to the FND’s environment. The concentrations were incrementally adjusted, and measurements were performed after each addition to evaluate the changes in the NV center properties.
II.4.2 pH Sensing
For the pH sensing experiments, an initial 0.1 M NaOH(aq) solution at pH 13 was prepared. The pH was then progressively decreased by adding controlled volumes of a 1 M HCl solution. After each pH adjustment, the properties of the NV centers were measured while the FNDs were maintained in the optical trap at 60 mW.
II.4.3 Temperature Sensing
For the temperature sensing experiments, the medium containing the FNDs was initially maintained at 37 C using an incubator. The temperature was then increased in steps of 1K, allowing the evaluation of temperature changes on the NV center properties. After each change of the set temperature of the incubator, we waited 10 minutes before the measurements were conducted.
III Results
We begin by presenting the main results of our experiments, starting with the examination of the general fluorescence properties of NV-centers within the FNDs through Photoluminescence (PL) spectroscopy. This foundational assessment provides an overview of the baseline optical characteristics and helps identify any immediate changes induced by the NIR laser irradiation. Following the PL spectrum analysis, we delve into the variations in fluorescence intensity as a function of optically detected magnetic resonance (ODMR) conditions. This step is crucial for understanding the impact of NIR irradiation on the magnetic resonance properties of the NV centers. Finally, we conduct T1-relaxometry to investigate the spin relaxation dynamics, providing a detailed examination of the spin-state populations and their evolution under different NIR laser parameters. This approach from general fluorescence to specific spin dynamics allows for a comprehensive evaluation of the NIR irradiation effects on the NV centers. PL spectrum measurement and ODMR measurement were conducted with CW NIR laser irradiation.
III.1 Photoluminescence Spectrum Analysis
We conducted PL spectroscopy of FNDs with varying NIR power. The FNDs were either dried on glass substrates or trapped in water to examine the influence of the surrounding medium.
Effect of NIR Laser Power on PL Spectrum Intensity
We exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW and for each power observe the PL spectrum from 550 nm to 750 nm with our spectrometer.
Figure 4: Effect of CW NIR laser on dry FND emission spectrum. Optical emission spectrum of dry FND on glass under CW NIR laser exposure at 0 mW (green), 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
Our PL measurements revealed significant alterations in the spectral intensities of both zero-phonon lines (ZPL) from the two NV charge states upon CW NIR laser exposure. Notably, as a function of NIR laser power there was a slightly more pronounced reduction in the spectral intensity of NV- compared to NV0. As shown in Fig. 4, the ZPL intensity of NV0 at 575 nm decreased by 29.5% and 32.4%, while the NV- ZPL intensity at 637 nm decreased by 52.8% and 62.3% when the CW NIR laser power was increased from 0 mW to 30 mW and 60 mW, for the condition with FNDs dry on glass. For the fluorescence spectrum (like Fig. 4), the curves remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30mW to 60mW (data not shown) and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water, cf. Fig. 5. Quantitatively, we observe a 1.2% general intensity difference from 30 mW to 60 mW for FNDs on glass with water while the same figure is 0.7% for FNDs trapped in water.
Figure 5: Effect of CW NIR laser on trapped FND emission spectrum. Optical emission spectrum of FND trapped in water CW NIR laser at 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
To determine the ratio of NV0 to NV- emission intensities under CW NIR laser powers of 0 mW, 30 mW, and 60 mW we integrated the emission intensities over the respective spectral ranges for NV0 (550–620 nm) and NV- (655–750 nm), but compared to the procedure of Ref. [13], we applied a normalization factor to account for the difference in spectral ranges:
$\frac{\text{NV}^{0}}{\text{NV}^{-}}=\left(\frac{\sum_{\lambda=550}^{620}I_{\lambda}}{\sum_{\lambda=655}^{750}I_{\lambda}}\right)\times\frac{95}{70}$ (1)
We observe that the NV0/NV- emission ratio stays the same for trapped FNDs, 0.45 for both 30mW and 60mW NIR laser power. In contrast, it rises from 0.39 without NIR irradiation to 0.51 with 60 mW NIR laser power for dry FND on glass. All results are displayed in Fig. 10.
III.2 Optically Detected Magnetic Resonance Measurements
We further examined the influence of CW NIR laser power on the ODMR resonance frequency and contrast for FNDs dried on glass substrates and those trapped in water.
Effect of CW NIR Laser Power on the Zero Field Splitting of the NV Center
We exposed the FNDs to CW NIR laser powers of 0 mW, 30 mW, and 60 mW, and measured the resulting ODMR spectra. Our primary observation was a shift in the resonance frequency with increasing CW NIR laser power, as illustrated with an example in Fig. 6.
Figure 6: Effect of CW NIR laser power on ODMR spectrum of dry FND. ODMR spectra of dry FNDs under CW NIR laser exposure at 0 mW (green) and 60 mW (red) of 1064 nm laser power. The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed.
For FNDs dried on glass substrates, the resonance frequency shifted significantly to the lower frequencies. Specifically, the resonance frequency shift increased in absolute value from $-0.31\pm 0.54$ MHz at 0 mW to $-5.83\pm 046$ MHz at 60 mW of CW NIR power; both values given relative to the expected 2.87 GHz. In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $-0.84\pm 0.21$ MHz at 30 mW and $-1.58\pm 0.33$ MHz at 60 mW (relative to the expected 2.87 GHz). Examples are shown in Fig. 6 and Fig. 7 while combined results are displayed in Fig. 11. As before, for the similar experiment when water was added to the same FND that remained stuck on the glass slide, the ODMR spectra show a trend very similar to Fig. 6 although the difference between 30 mW and 60 mW of NIR power is less pronounced (data not shown).
Figure 7: Effect of CW NIR laser power on ODMR spectrum of trapped FND. ODMR spectra of FNDs trapped in water under CW NIR laser exposure at 30 mW (orange) and 60 mW (red) of 1064 nm laser power. The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed.
III.3 T1 Relaxometry Measurements
We conducted $T_{1}$ relaxometry experiments under different NIR laser power. As above, FNDs were either dried on glass substrates or suspended in water to assess the influence of the surrounding medium on the T1 relaxometry measurements with NIR irradiation. The samples were exposed to NIR laser powers of 0 mW, 30 mW, and 60 mW, and the corresponding $T_{1}$ values were measured. For these experiments, the pulsing sequence in Fig. 3 was applied.
Effect of NIR Laser Power on $T_{1}$ Relaxation Times
First, we investigate how the T1 value of dry FNDs is affected by NIR irradiation. For FNDs dried on glass substrates, we observed a significant decrease in the $T_{1}$ relaxation times with increasing NIR laser power. As shown in the example in Fig 8, the mean $T_{1}$ value decreased from $241\pm 21~{}\mu$s at 0 mW NIR power to $136\pm 19~{}\mu$s at 60 mW NIR power. When water was added to the same FND that remained stuck on the glass slide, the $T_{1}$ values showed very similar changes. The mean $T_{1}$ value decreased from $237\pm 25~{}\mu$s at 0 mW NIR power to $154\pm 17~{}\mu$s at 30 mW NIR power and $143\pm 21~{}\mu$s at 60 mW NIR power. (data not shown).
Figure 8: Effect of NIR laser power on $T_{1}$ relaxation times of dry FNDs. $T_{1}$ relaxometry measurements for FNDs dried on glass under NIR laser powers of 0 mW (green) and 60 mW (red) at 1064 nm wavelength. Pulsing as in Fig 3.
For FNDs trapped in water, the decrease in $T_{1}$ due to NIR irradiation was significantly less. As depicted in the example in Fig. 9, we do not observe any significant change in the 2 curves. The $T_{1}$ values for trapped FNDs were $179\pm 8~{}\mu$s at 30 mW and $176\pm 8~{}\mu$s at 60 mW, showing a negligible decrease compared to the values at 0 mW, with all results summarized in Fig. 11.
Figure 9: Effect of NIR laser power on $T_{1}$ relaxation times of trapped FNDs. $T_{1}$ relaxometry measurements for FNDs trapped in water under pulsed NIR laser powers of 30 mW (orange) and 60 mW (red) at 1064 nm wavelength.
III.4 Combined Statistical Analysis
In this section, we present a comprehensive statistical analysis of our experimental results.
Relationship Between NV0/NV- Ratio and ODMR Frequency Shift
Figures 10.a, 10.c, 11.a and 11.b illustrate the dependence of the NV0/NV- fluorescence intensity ratio and the ODMR resonance frequency shift for FNDs under varying CW NIR laser powers.
For FNDs dried on glass substrates, we observe a significant increase in the NV0/NV- fluorescence intensity ratio and a corresponding shift in the ODMR resonance frequency with increasing CW NIR power from 0 mW to 60 mW.
In contrast, for FNDs trapped in water, the NV0/NV- fluorescence intensity ratio and ODMR frequency shift exhibit much smaller changes with increasing CW NIR power. This suggests that the aqueous environment mitigates the effects of CW NIR irradiation.
Correlation Between $T_{1}$ Relaxation Times and ODMR Frequency Shift
Figures 10.b, 10.d, 11.a and 11.c show the relationship between the $T_{1}$ relaxation times and the ODMR resonance frequency shift for FNDs under varying NIR laser powers. $T_{1}$ was obtained with the pulse sequence as in Fig 3 while the ODMR was obtained with CW NIR irradiation.
For FNDs dried on glass substrates, there is an effect between $T_{1}$ times and ODMR frequency shifts with increasing NIR power. The significant decrease in $T_{1}$ times and increase in frequency shift suggest that NIR irradiation enhances spin relaxation processes.
For FNDs trapped in water, the $T_{1}$ relaxation times remain relatively constant, and the ODMR frequency shifts exhibit minimal changes with increasing NIR power.
Overall, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts. Further, our results confirm that with specific NIR irradiation sequences and optimizing laser parameters, cf. Fig. 3 and Ref. [19], potential detrimental effects of NIR on $T_{1}$ relaxation can be significantly mitigated, even with FNDs dry on glass.
Figure 10: a) Correlation between NIR laser power, ODMR resonance frequency shift and NV0/NV- fluorescence intensity ratio for FNDs trapped inside water. b) Correlation between NIR laser power, ODMR resonance frequency shift and T1-relaxometry times for FNDs trapped inside water. c) Correlation between NIR laser power, ODMR resonance frequency shift and NV0/NV- fluorescence intensity ratio for FNDs dry on glass. d) Correlation between NIR laser power, ODMR resonance frequency shift and T1-relaxometry times for FNDs dry on glass. Data for 30 FNDs exposed to 0 mW (green), 30 mW (orange), and 60 mW (red) of 1064 nm NIR laser irradiation. CW NIR for ODMR and spectra, pulsed NIR as in Fig 3 for T1-relaxometry. In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and no such trend for FNDs in the optical trap in water.
Figure 11: Summary of experimental results, showing variation for individual nanodiamonds. (a) ODMR resonance frequency shift, (b) NV0/NV- fluorescence intensity ratio, and (c) $T_{1}$ relaxometry times as functions of NIR laser power. Data for 5 random FNDs dry on glass (black stars) and trapped in water (gray circles) with 1064 nm NIR laser irradiation.
Statistical Analysis Using ANOVA
To study the significance of the observed effects, we performed an Analysis of Variance (ANOVA) test on the $T_{1}$ relaxation times, ODMR frequency shifts, and NV0/NV- fluorescence intensity ratios. The ANOVA results are summarized in Table 1.
Table 1: ANOVA test results for the effect of NIR laser power on FNDs dried on glass substrates, on glass after adding water and trapped in water. The p-values indicate the significance of differences observed across NIR power levels for $T_{1}$ relaxation times, ODMR frequency shifts, and NV0/NV- ratios.
p-value
FNDs dried on glass
$T_{1}$ Relaxation Time $<0.0001$
ODMR Frequency Shift $<0.0001$
NV0/NV- Ratio 0.0024
FNDs on glass with water
$T_{1}$ Relaxation Time $0.072$
ODMR Frequency Shift $0.008$
NV0/NV- Ratio 0.201
FNDs Trapped in Water
$T_{1}$ Relaxation Time 0.221
ODMR Frequency Shift 0.073
NV0/NV- Ratio 0.371
For FNDs dried on glass substrates, the very low p-values (p $<0.01$) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant. This indicates that NIR irradiation has a considerable impact on the NV centers’ properties in air.
For FNDs on glass substrates after adding water, the medium p-values ($0.20.01<p<$ 0.2) for all three parameters confirm that the differences observed with varying NIR laser powers are not always statistically significant. In fact, the difference between no NIR laser and NIR laser is significant but the differences between 30 mW and 60 mW power of NIR laser are not significant.
In contrast, for FNDs trapped in water, the p-values are high (p $>0.07$), suggesting that the differences observed across different NIR power levels are not statistically significant. This reinforces the conclusion that the aqueous environment mitigates the effects of NIR irradiation on the NV centers.
III.5 Charge State Dynamics
To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model [32]. This model accounts for the various optical transitions, inter-system crossings, ionization, and recombination processes that occur when NV centers are simultaneously exposed to green and NIR laser excitation.
The NV center can exist in two charge states: the negatively charged NV- and the neutrally charged NV0. Each charge state has its own ground and excited electronic states, as well as inter-system crossing pathways to metastable singlet states. The model considers 8 states as illustrated in the schematic energy level diagram in Fig. 12.
The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state [32]. The rates of change had to be adapted to our experimentation with green 532 nm and NIR 1064 nm laser.
The model is based on the following assumptions:
  • Steady-State Approximation: The populations reach a steady state within the simulation time, allowing us to analyze the equilibrium behavior.
  • Neglect of Higher-Order Processes: Multi-photon processes and higher excited states are neglected due to their low probabilities under the experimental conditions.
  • Constant Rate Coefficients: The rate coefficients are assumed to be constant and independent of the local temperature and environmental factors.
  • No External Fields: Effects of external magnetic or electric fields on the transition rates are not included.
Figure 12: Energy level diagram of the NV center, showing the ground states (P1, P2 and P7), excited states (P3, P4 and P8), singlet states (P5 and P6), and the transitions between them. Solid arrows represent radiative transitions, dotted arrows represent non-radiative transitions, and dashed arrows represent ionization and recombination processes.
Population Dynamics of NV- and NV0 Centers
The simulation results reveal that at low CW NIR laser powers, the population of NV- centers increases slightly, while the NV0 population decreases. This is attributed to enhanced recombination processes facilitated by the NIR photons, which convert NV0 centers back to NV-. Specifically, up to an CW NIR power of approximately 10 mW, the NV- population increases by about 8% compared to the case without CW NIR illumination.
However, as the CW NIR power increases beyond this threshold, the trend reverses. The NV- population begins to decrease, and the NV0 population increases. At a CW NIR power of 60 mW, the NV- population decreases by approximately 10% from its maximum value, while the NV0 population increases correspondingly. This inversion is due to the saturation of recombination processes and the dominance of ionization processes at higher CW NIR powers, which ionize NV- centers back to NV0. The threshold depends on the recombination rate, ionization rate and on the initial population of each state.
Fluorescence Intensity and NV0/NV- Ratio
The total fluorescence intensity from the NV centers decreases with increasing CW NIR power (Fig. 13). This decrease is primarily due to the reduced population of NV- centers and the lower fluorescence quantum yield of NV0 centers. The fluorescence from NV- centers dominates the emission under low CW NIR power, but as the CW NIR power increases, the contribution from NV0 centers becomes more significant and we observe a saturation in the fluorescence.
The ratio of NV0 to NV- fluorescence intensities ($F_{\text{NV}^{0}}/F_{\text{NV}^{-}}$) increases with CW NIR power. At low CW NIR powers, this ratio is low due to the predominance of NV- fluorescence. As the CW NIR power increases, the ratio rises slowly, indicating a relative increase in NV0 emission. This change in the fluorescence ratio affects the overall spectral characteristics and can influence the detection sensitivity in applications relying on NV- fluorescence.
Consistency with Experimental Observations
The trends observed in the simulation align with the experimental results. The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity and NV0/NV- fluorescence intensity ratio observed experimentally. The simulation provides a quantitative framework to understand how NIR illumination influences the charge state dynamics and, consequently, the optical and spin properties of NV centers.
Figure 13: Total fluorescence intensity from NV centers as a function of NIR laser power. (Black: Theory, Red: Experimental observations.
III.6 Temperature Simulation
To model the temperature increase induced by the NIR laser on FNDs, we simulated the heat transfer. The simulation accounts for the absorption of laser energy by the FND and the subsequent heat dissipation into the surrounding medium (water or air). The key processes considered are:
  1. 1.
    Energy Absorption: The FND absorbs a portion of the incident laser power, leading to an increase in its internal energy and temperature.
  2. 2.
    Heat Loss: Heat is lost from the FND to the surrounding medium through conduction and convection, depending on the thermal properties of the medium.
The simulation is based on the following assumptions:
  • Uniform Temperature Distribution: The FND is assumed to have a uniform temperature at any given time due to its small size.
  • Negligible Radiation Losses: Heat loss due to thermal radiation is neglected because it is insignificant compared to conduction at the temperatures considered.
  • Constant Thermal Properties: The thermal properties of the FND and the medium are assumed constant over the temperature range studied.
  • Steady-State Heat Transfer Coefficient: The convective heat transfer coefficient $h$ is considered constant and based on the thermal conductivity of the medium.
  • No Phase Change: The medium remains in the same phase (liquid for water, gas for air) throughout the simulation.
Temperature Evolution During a 500 ns NIR Pulse
[…]
In Water: The temperature increase of the FND is minimal due to the high thermal conductivity and heat capacity of water. The simulation indicates that during the 500 ns NIR pulse:
  • The maximum temperature rise ($\Delta T_{\text{max}}$) is approximately 2 K above the ambient temperature.
  • The temperature increases steadily during the laser pulse and rapidly returns to ambient temperature within nanoseconds after the laser is turned off.
In Air: The temperature increase is more pronounced because air has lower thermal conductivity and heat capacity. The simulation results show:
  • The maximum temperature rise reaches approximately 41 K above the ambient temperature during the 500 ns NIR pulse.
  • The temperature increases during the laser pulse and decreases more slowly than in water after the laser is turned off.
Consistency with Experimental Observations
The simulation results align with experimental observations where pulsed NIR illumination causes negligible thermal effects in aqueous environments. This supports the conclusion that charge state dynamics are the primary mechanism affecting NV center behavior under short NIR pulses in water.
In experiments conducted in air, temperature increases contribute to shifts in ODMR frequency and should be accounted for. These shifts correspond to a 2 MHz shift in the ODMR resonance frequency which is what we experimentally observe.
IV Discussion: Effects of NIR Light on FNDs
Our experimental and simulation results reveal a complex interplay between thermal dynamics and charge state transitions in NV within FNDs under NIR illumination. The observed effects include a decrease in total fluorescence intensity, reduction in spin-lattice relaxation time, shift in the ODMR resonance frequency, decrease in ODMR contrast, and an increase in the NV0/NV- fluorescence ratio.
IV.1 Thermal Effects Induced by NIR Illumination
One primary mechanism contributing to these observations is the localized heating induced by the NIR laser. The shift in the ODMR resonance frequency with increasing NIR power, as shown in Figures 6, 7 and 10, and 11a), suggests an increase in temperature within the FNDs. This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately $-74$ kHz/K [33, 34].
However, the extent of the thermal effect is significantly influenced by the surrounding medium’s thermal properties. Media with higher thermal conductivity and heat capacity, such as water, dissipate heat more efficiently than air, resulting in smaller temperature increases in the FNDs. This is supported by our temperature simulations, which show minimal temperature rise in water compared to air during NIR illumination. Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts and negligible changes in $T_{1}$ times with increasing NIR power, as shown in Figures 7, 10 and 11.
IV.2 Charge State Dynamics and Fluorescence Changes
In addition to thermal effects, the NIR laser alters the charge state dynamics of the NV centers. Our simulations of charge state dynamics indicate that at low NIR power levels, there is a slight increase in the NV- population due to enhanced recombination processes facilitated by NIR photons. However, as the NIR power increases beyond a threshold (approximately 10 mW with our parameters), the NV- population begins to decrease while the NV0 population increases. This inversion is attributed to the saturation of recombination processes and increased ionization rates at higher NIR powers.
These changes in charge state populations directly affect the fluorescence properties of the FNDs. The total fluorescence intensity decreases with increasing NIR power, as shown in Figure 13. This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers [35]. The increase in the NV0/NV- fluorescence ratio with higher NIR power, depicted in Figures 10 and 11, further confirms the shift in charge state distribution.
Additionally, the increased NV0 background fluorescence dilutes the ODMR contrast since NV0 centers do not exhibit spin-dependent fluorescence variations. This is reflected in the decreased ODMR contrast observed in our measurements (Figures 6-7). Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{1}$ times [20, 36].
IV.3 Interplay Between Thermal Effects and Charge State Dynamics
Although both thermal effects and charge state dynamics significantly impact the properties of the NV centers under NIR illumination, their relative contributions depend on the environmental conditions. In FNDs dried on glass substrates, the pronounced ODMR frequency shifts suggest that thermal effects dominate due to inefficient heat dissipation in air.
In contrast, for FNDs trapped in water, the minimal temperature increase reduces thermal effects, making charge-state dynamics more prominent in influencing the NV centers’ behavior. The increase in the NV0/NV- ratio and the decrease of T1-relaxometry times with NIR power without significant ODMR frequency shifts (Figure 10) indicates that charge state transitions are the primary mechanism affecting the fluorescence and spin properties in aqueous environments.
IV.4 Alternative Explanations and Considerations
Although thermal effects and charge state dynamics provide a comprehensive explanation for our observations, alternative mechanisms should be considered. Multi-photon absorption processes could directly excite the NV centers without significant heating. However, the relationship between ODMR frequency shift and NIR power (Figures 7 and 10) supports the thermal interpretation in our experiments.
Moreover, the significant differences in the effects of NIR irradiation between FNDs in air and water emphasize the critical role of the surrounding medium. The statistical analysis using ANOVA (Table 1) confirms that the differences observed in FNDs dried on glass are statistically significant, whereas those in FNDs trapped in water are not, reinforcing the influence of environmental thermal properties.
V Sensing Capabilities of Trapped FNDs
As a final step in the comprehensive investigations, this section explores the potential of optically trapped FNDs with NV centers for sensing applications. Experiments focused on three key parameters: Gd3+ concentration, pH, and temperature. Measurements included ODMR, T1 relaxation, and optical spectrum analysis under variation of the three parameters mentioned above.
Figure 14: Variation in three sensing parameters when subject to controlled variation of paramagnetic species, pH, and temperature. a) ODMR resonance frequency shift VS Gd3+ concentration. b) ODMR resonance frequency shift VS pH value. c) ODMR resonance frequency shift VS Temperature increase. d) T1-relaxometry time VS Gd3+ concentration. e) T1-relaxometry time VS pH value. f) T1-relaxometry time VS Temperature increase. g) NV0/NV- fluorescence intensity ratio VS Gd3+ concentration. h) NV0/NV- fluorescence intensity ratio VS pH value. i) NV0/NV- fluorescence intensity ratio VS Temperature increase. The boxplot represent results for 10 FNDs that were trapped for each condition with 60 mW 1064 nm CW NIR laser. In each case, results for Gd3+ 0M, pH 13 and 0K trapped were used as reference values for the individual FNDs. The red squares emphasize the optimal measurement method for each sensing parameter. The ODMR resonance frequency value is given relative to the expected 2.87 GHz.
V.1 pH Sensing
The impact of environmental pH on the properties of NV centers was evaluated using buffer solutions with pH values of 1, 7, and 13.
Figure 15: NV0/NV- fluorescence intensity ratio during pH value changes. Measurement of the same FND while changing from pH13 to pH1 by adding HCl solution (Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
  • T1 Relaxation: A decrease in T1 values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig. 14.h). This is consistent with reports that high pH conditions stabilize spin relaxation dynamics [1].
  • Optical Spectrum: A decrease in the NV0/NV- fluorescence intensity ratio with increasing pH (Fig. 14.e and  15) suggests that alkaline conditions favor the NV- charge state, possibly due to deprotonation of surface functional groups, which enhances negative surface potential [37]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser shows a clear advantage of using FNDs in suspension for sensing.
  • ODMR: No significant changes in resonance frequency (Fig. 14.b) were detected across the tested pH range, indicating minimal impact of pH on electronic spin states.
The NV0/NV- fluorescence intensity ratio’s dependence on pH underscores the importance of surface chemistry in modulating charge state dynamics, especially with FNDs in suspension.
V.2 Paramagnetic Species Sensing
The sensitivity of NV centers to paramagnetic gadolinium ions (Gd3+) was investigated through measurements at 3 different values of the concentration.
Figure 16: NV0/NV- fluorescence intensity ratio during gadolinium concentration changes. Measurement of the same FND while changing gadolinium concentration from 0M to 0.1M by adding a gadolinium solution (0.45 M in acetate) (Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
  • T1 Relaxation: A significant decrease in T1 values was observed with increasing Gd3+ concentration (Fig. 14.g). This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment [38].
  • Optical Spectrum: A decrease in the NV0/NV- fluorescence intensity ratio was observed (Fig. 14.d and  16). This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 [39]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser shows a clear advantage of using FNDs in suspension for sensing.
  • ODMR: No significant shifts or changes in resonance frequency were detected, suggesting limited interaction between Gd3+ ions and the electronic spin states of the NV centers under the tested conditions (Fig. 14.a).
These findings demonstrate the potential of NV centers for detecting paramagnetic species like Gd3+. The sensitivity saturates at higher concentrations due to the finite interaction volume and saturation of surface sites for Gd3+ binding. The T1-relaxometry provides a robust method for quantifying Gd3+ in biological and environmental contexts. However, the effect on charge state dynamics also imply that conclusions from T1-relaxometry in otherwise uncharacterized environments should be made with caution, and at best complemented with fluorescence spectrometry. For completeness, we observe that by adding Gd3+ to a concentration of 0.1 M, the pH value in the solution increases slightly (by of order 1-2 pH units) which is consistent with our previous observations on the effect of pH on the fluorescence of FNDs..
V.3 Temperature Sensing
The temperature response of NV centers was evaluated under three conditions: FNDs on glass substrates with no NIR laser irradiation, trapped FNDs with no temperature increase, and trapped FNDs with temperature increases of +1 K and +2 K.
[…]
  • T1 Relaxation: No significant changes were detected in T1 values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig. 14.i).
  • Optical Spectrum: No significant changes in the NV0/NV- fluorescence intensity ratio were observed with temperature, suggesting that charge state dynamics are minimally affected by thermal fluctuations (Fig. 14.f).
ODMR’s temperature sensitivity highlights its utility for high-resolution nanoscale thermometry. The absence of significant T1 or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements
VI Conclusion
This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive understanding of the interplay between thermal effects, charge state dynamics, and their combined influence on sensing capabilities. By employing an integrated experimental approach combining photoluminescence spectroscopy, optically detected magnetic resonance (ODMR), and T1 relaxometry, this study elucidates how varying NIR laser powers impact the fluorescence, spin properties, and stability of NV centers in diverse environmental conditions.
A key finding is the dual role of NIR irradiation: in dry environments, thermal effects dominate, resulting in significant shifts in ODMR frequencies and reductions in T1 relaxation times due to localized heating. In contrast, for FNDs that are fully submerged in aqueous environments, the surrounding water appears to mitigate these thermal effects through efficient heat dissipation, allowing charge state dynamics to take precedence. The increased NV-, NV0 fluorescence intensity ratio and its implications for sensing applications underscore the critical role of the surrounding medium in tailoring NV center behavior. The study also demonstrates the sensitivity of optically trapped FNDs to pH variations, temperature fluctuations, and paramagnetic species, establishing these nanodiamonds as versatile tools for nanoscale sensing.
The results have profound implications for the development of robust NV-based sensing protocols. The findings indicate that optimizing NIR laser parameters, including power levels and pulsing sequences, can significantly enhance sensor performance while minimizing adverse effects such as fluorescence quenching and spin relaxation. Additionally, the demonstrated ability to detect and differentiate environmental factors such as pH and gadolinium ions highlights the potential of NV centers for applications in biological sensing and environmental monitoring.
However, this study also reveals critical uncertainties and challenges that must be addressed to fully exploit the potential of NV-based FNDs. Long-term stability under extended NIR irradiation remains an open question, particularly regarding potential degradation of fluorescence and spin properties. The effects of complex biological environments, including interactions with proteins, ions, and other cellular components, are yet to be fully understood. The role of surface functionalization in enhancing NV center stability and sensing accuracy requires further exploration, as does the potential for alternative NIR wavelengths or advanced laser modulation techniques to improve performance.
Theoretical modeling, while providing valuable insights into charge state dynamics and thermal effects, must be expanded to include the complexities of real-world applications. Future work could explore the integration of complementary sensing techniques, such as fluorescence lifetime imaging or advanced magnetic resonance methods, to provide a more holistic understanding of NV center behavior under varying conditions. Furthermore, efforts to miniaturize and automate the experimental setup, enabling in vivo applications, would be a critical step toward translating this research into practical diagnostic tools.
In conclusion, this study lays a strong foundation for the continued development of NV-based sensing technologies, bridging fundamental insights with application-driven innovations. By addressing the remaining uncertainties and refining experimental protocols, FNDs with NV centers have the potential to become transformative tools in nanoscale sensing, advancing fields ranging from cellular biology to nanoscale physics.
Acknowledgements.
This work was supported by the Independent Research Fund Denmark (grant no. 0135-00142B) and the Novo Nordisk Foundation (grant no. NNF20OC0061673). We acknowledge discussions with Maabur Sow and Fedor Jelezko, supported by ERC through HyperQ Project (SyG 856432). We further acknowledge insights provided through the student project work of Kristian Lambertsen and Nicolai Kongstad, including the sketch of the experimental setup, Fig. 2.
appendix
As a varying range of values for the rate constants used in the numerical model exist in literature [23, 32], we provide the values used in the simulations presented here.
Table 2: Parameters used in the simulation
Parameter Value Units Description
$K_{\mathrm{NV}^{-}}^{(e)}$ 27 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ NV- excitation rate
$K_{\mathrm{NV}^{0}}^{(e)}$ 18 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ NV0 excitation rate
$\tau_{\mathrm{NV}^{-}}$ $12\times 10^{-3}$ $\mu\mathrm{s}$ Excited-state lifetime (12 ns)
$\tau_{\mathrm{NV}^{0}}$ $20\times 10^{-3}$ $\mu\mathrm{s}$ Excited-state lifetime (20 ns)
$K_{\mathrm{NV}^{-}}^{(f)}$ $1/\tau_{\mathrm{NV}^{-}}$ $(\mu\mathrm{s})^{-1}$ NV- fluorescence rate
$K_{\mathrm{NV}^{0}}^{(f)}$ $1/\tau_{\mathrm{NV}^{0}}$ $(\mu\mathrm{s})^{-1}$ NV0 fluorescence rate
$K_{35}$ 7.9 $(\mu\mathrm{s})^{-1}$ NV- ES ms=0 $\to$ singlet
$K_{45}$ 45 $(\mu\mathrm{s})^{-1}$ NV- ES ms=±$\pm 1$to singlet
$K_{56}$ 1000 $(\mu\mathrm{s})^{-1}$ Singlet ES $\to$ GS
$K_{61}$ 6.5 $(\mu\mathrm{s})^{-1}$ Singlet GS $\to$ NV- GS ms=0
$K_{62}$ 0.1 $(\mu\mathrm{s})^{-1}$ Singlet GS $\to$ NV- GS ms=±$\pm 1$
$K_{iG}$ 852 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Ionization rate (green laser)
$K_{rG}$ 134 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Recombination rate (green laser)
$K_{iIR}$ 1.2 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Ionization rate (IR laser)
$K_{rIR}$ 3.17 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Recombination rate (IR laser)
laser_green_power 0.7 mW Green laser power

### v2

@@
 Biosensing based on optically trapped fluorescent nanodiamonds is an intriguing research direction potentially allowing to resolve biochemical processes inside living cells.
-Towards this goal, we investigate infrared near (NIR) laser irradiation at 1064 nm on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers.
+Towards this goal, we investigate near infrared (NIR) laser irradiation at 1064 nm on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers.
+The NIR wavelength is a popular choice for optical trapping in biological samples because of its low absorption in bio-samples.
 By conducting comprehensive experiments, we aim to understand how NIR exposure influences the fluorescence and sensing properties of FNDs and to determine the potential implications for the use of FNDs in various sensing applications.
 The experimental setup involved exposing FNDs to varying intensities of NIR laser light and analyzing the resultant changes in their optical and physical properties.
-Key measurements included T1 relaxation times, optical spectroscopy, and optically detected magnetic resonance (ODMR) spectra.
+Key measurements included all-optical fluorescence relaxation, optical spectroscopy, and optically detected magnetic resonance (ODMR) spectra.
 The findings reveal how increased NIR laser power correlates with alterations in ODMR central frequency but also that charge state dynamics under NIR irradiation of NV centers plays a role.
@@
 preprint: APS/123-QED
-Figure 1: Graphical abstract illustrating the exploration of effects of NIR laser on NV center for bio-sensing.
 I Introduction
-Fluorescent nanodiamonds (FNDs) containing Nitrogen-Vacancy (NV) centers have attracted significant interest due to their combined optical and magnetic properties, which render them highly versatile for applications in particular in biomedical sensing and imaging [1, 2].
-The capacity of NV centers for optical polarization and readout of spin states earmarks them as formidable tools for advanced sensing applications [3].
-Yet, the utility of these sensors is heavily influenced by the photodynamics of the NV centers’ spin- and charge-states, which in turn governs their sensitivity and accuracy [4].
-NV centers possess an electronic spin S=1 system, characterized by spin sublevels $m_{s}=0$ and $m_{s}=\pm 1$, which are separated by a zero-field splitting due to spin-spin interaction.
-This zero-field splitting is sensitive to external parameters such as temperature, strain, and electric fields, while the energies of the spin sublevels can be further shifted by magnetic fields through the Zeeman effect [5, 6].
-Additionally, the optical spectrum of NV centers, including their fluorescence intensity and wavelength, is determined by the charge state distribution and depends on the local electrochemical potential and intrinsic charge state dynamics.
-These dependencies allow NV centers to act as versatile nanoscale sensors, providing insights into local magnetic fields across various frequency ranges, temperature fluctuations, and surface electrochemical conditions.
-Therefore, to accurately measure external parameters like magnetic fields or temperature with high sensitivity in biological environments, it is essential to understand how external control—such as a near infrared trapping laser—affect the intrinsic properties of NV centers [7, 8].
-Trapping of FNDs with optical tweezers is an attractive approach as it potentially allows for spatially resolved sensing in challenging environments, with a resolution determined by the size of the FND and of the trapping potential [9].
-To achieve this in biological samples and minimize cell and tissue damage, the wavelength of the intense trapping laser must be in the biological transparency window, such as 1064 nm, where residual absorption of tissue components is minimal [10].
-In this work, we employed three principal methods to investigate the properties of NV centers in FNDs under NIR light exposure: spin polarization relaxometry (T1-relaxometry) [11, 12], optically detected magnetic resonance (ODMR) spectroscopy [13, 14], and photoluminescence (PL) spectrum analysis [15].
-PL spectroscopy involves measuring the emitted light from the NV centers when excited by a laser, providing information on the optical properties and charge states of the NV centers [16].
-ODMR combines optical and microwave techniques to detect changes in the NV centers’ fluorescence in response to varying magnetic fields and relative temperature changes, allowing for precise measurement of the NV centers’ electron spin states and local magnetic environment [17].
-T1-relaxometry, on the other hand, measures the longitudinal relaxation time (T1), which is the time it takes for the NV centers’ spin states to return to thermal equilibrium after being polarized [18].
-This method is valuable for sensing processes and can detect various types of magnetic noise at the spin-transition frequency, such as free radicals [12].
-Combined, these techniques provide a comprehensive analysis of the NV centers’ optical, magnetic, and spin properties under varying experimental conditions [4].
-We employ a controlled 30 to 60mW NIR laser, with a pulsing sequence inspired by [19], and suggest that the approach can facilitate more precise and reliable intracellular biosensing assessments while keeping any undesirable NIR effect low.
-Previous efforts by Aslam et al.
-[20] have already explored the impact of near-infrared (NIR) laser irradiation on the properties of NV centers, revealing that NIR irradiation can significantly modify NV photoluminescence and enhance spin-lattice relaxation processes, using factors such as NIR laser power (around 30mW) and pulse duration (around 30ns) [21].
-The existing literature provides different and sometimes conflicting explanations on the physical mechanisms induced by NIR radiation, which modify both the optical and spin properties of NV centers [22, 15, 9, 21, 23, 14].
-In our experiment, we first investigated the effect of the NIR laser on the fluorescence of FNDs in two scenarios: dry FNDs and optically trapped FNDs.
-We then demonstrated the sensing capabilities of the optically trapped FNDs, which included measurements in different pH environments, controlled temperature variations, and the addition of a paramagnetic species (Gd3+).
-We propose that changes in spin and photoluminescence properties may be due to three effects: i) heating of the FNDs induced directly by NIR laser irradiation of the FNDs by the NIR laser, ii) a charge state recombination and ionization rate that depends on the NIR laser power and iii) a sensitivity effect, where the NIR laser alters the FNDs’ response to their environment [24].
-The sensitivity effect is particularly important for those sensing protocols that rely on NV centers positioned close to the surface of the FNDs [4, 17].
-For example, the pH of the surrounding environment plays a significant role in the behavior of NV centers in FNDs.
-Variations in pH can influence the charge state dynamics of NV centers, which in turn affects their fluorescence properties and overall sensitivity in sensing applications [24, 18, 13, 25].
-In this article, we build upon existing hypotheses that the effects of NIR irradiation on NV centers involve both heating, as suggested by previous studies [22], and redistribution between charge states [21, 23].
-We aim to provide a more comprehensive understanding of these combined effects, thus offering new insights into optimizing optically trapped NV-based sensing protocols [9].
-The paper is organized as follows: Section II describes the methods, including sample preparation, the optical setup, FND trapping, and sensing experiments.
-Section III presents the results, focusing on photoluminescence spectrum analysis, optically detected magnetic resonance measurements, T1 relaxometry, and a combined statistical analysis.
-Section IV discusses the implications of the findings, including thermal effects, charge state dynamics, and their interplay, along with alternative considerations.
-Finally, Section V explores the sensing capabilities of trapped FNDs.
+Figure 1: Graphical abstract illustrating the sensing modalities explored to investigate effects of NIR laser on NV center for sensing.
+Temperature, abundance of paramagnetic species, and pH, are altered while optical spectroscopy, ODMR spectra, and fluorescence polarization relaxometry is measured from optically trapped nanodiamonds.
+Fluorescent nanodiamonds (FNDs) hosting negatively-charged nitrogen-vacancy (NV-) centers combine bright, photostable fluorescence with quantum-coherent electron-spin states, making them uniquely suited for bio-sensing and bio-imaging at the nanoscale [Schirhagl2014, Liu2021].
+Because NV spins can be optically polarised and read out at room temperature, changes in their spin-state populations translate directly into optical signals, enabling quantitative sensing of magnetic fields, temperature and local electro-chemical potentials [Doherty2013, Rondin2014].
+The ultimate performance of those sensors, however, is governed by the spin- and charge-state photodynamics of the NV centers themselves; understanding how an experimental control parameter perturbs those dynamics is therefore a prerequisite for accurate, high-sensitivity measurements.
+The NV center is an electronic spin-1 system with a ground-state zero-field splitting (ZFS) $D\approx 2.87~\mathrm{GHz}$ that separates the $m_{s}=0$ and $m_{s}=\pm 1$ sub-levels.
+External magnetic fields lift the degeneracy of the $m_{s}=\pm 1$ levels through the Zeeman effect, while lattice strain, temperature and electric fields shift the ZFS and the optical transition energies via the Stark effect [Maze2008, Dolde2011].
+In addition, the relative abundance of the neutral (NV0) and negative (NV-) charge states—and thus the fluorescence spectrum and intensity—depends sensitively on the local Fermi level and on photo-ionisation/recombination pathways.
+These intertwined dependencies let NV centers act as multi-modal nano-probes, provided that the separate interaction between all external controls (laser beams, microwave fields, etc.) and the FNDs are themselves well understood and do not compromise sensor performance.
+A particularly powerful external control is optical trapping.
+Near-infra-red (NIR) optical tweezers can confine—and manoeuvre—sub-100  nm diamonds inside living cells or complex micro-fluidic environments, offering spatially resolved sensing with a resolution set by the nanoparticle size rather than by the optical diffraction limit [Neukirch2013].
+Operating the trap at 1064  nm keeps tissue absorption low and mitigates photo-damage [Svoboda1994], yet the same intense NIR field may also heat the nanodiamond, alter NV charge-state kinetics or modify spin-lattice relaxation.
+Earlier work on bulk diamond or levitated nanodiamonds indeed reported substantial NIR-induced changes in fluorescence and $T_{\rm relax}$[Aslam2013, Ji2016], but a clear picture for FNDs dispersed in aqueous media—arguably the configuration of greatest practical relevance-is missing.
+Here, as illustrated by Fig.1, we address that gap.
+Using freely diffusing and optically trapped FNDs in water and in biologically relevant buffer solutions, we quantify how a 1064  nm focused trapping beam (30–60  mW average power) affects NV optical and spin properties.
+We employ tree complementary read-outs:
+  1. 1.
+    photoluminescence (PL) spectroscopy to monitor charge-state redistribution [Geiselmann2013];
+  2. 2.
+    optically detected magnetic resonance (ODMR) to track shifts in the resonance frequencies with kHz precision [Gruber1997, Sow2020]; and
+  3. 3.
+    fluorescence relaxometry to probe magnetic-noise spectra—e.g. from free radicals or $\mathrm{Gd}^{3+}$ ions—at the NV transition frequency [Hall2016].
+Together, these observables allow us to disentangle three mechanisms that NIR light can induce:
+  1. 1.
+    Photothermal heating of the FND, shifting the ZFS via the temperature coefficient $dD/dT\approx-74~\mathrm{kHz\,K^{-1}}$[Acosta2010];
+  2. 2.
+    Power-dependent ionisation/recombination that re-weights NV- and NV0, thereby changing both ODMR spectral properties and PL brightness [Aslam2013, Ji2016];
+  3. 3.
+    Modified surface sensitivity, whereby the NIR field—through either heating or altered charge state—amplifies or suppresses the NV response to environmental parameters such as pH or paramagnetic species concentration [Fu2010].
+We first benchmark these effects in air-dried FNDs, then in trapped FNDs immersed in solutions of controlled pH and temperature, and finally in the presence of a paramagnetic relaxant ($\mathrm{Gd}^{3+}$).
+By comparing experiments with and without the NIR trap, we establish quantitative limits on NIR-induced artefacts and delineate operating regimes where accurate, spatially resolved biosensing is possible.
 II Methods
@@
 Fluorescent FNDs with a nominal diameter of 120 nm (Sigma-Aldrich, 798088) were used for most of the experiments reported in this article if not stated otherwise.
-They possess a significant number of nitrogen-vacancy (NV) centers due to high concentration of 3ppm.
+They possess a significant number of nitrogen-vacancy (NV) centers due to a high concentration of 3ppm.
 To prepare the FND suspension, 1 µL of the commercial stock solution of FNDs was diluted in 300-700 µL of deionized water depending on the desired concentration and before sonicating for 10 minutes.
 Next, the FND suspension was deposited on a petri dish (Mattek, part no P35G-1.5-10-C.), thoroughly cleaned to remove any contaminants that could interfere with the experiments and plasma ionised in order to make the surface hydrophilic, ensuring an even distribution of FNDs on the surface.
-Plasma treatment was applied for 30 seconds in a simple custom-device with a closed chamber in PMMA and under ambient conditions in which an electrode (Electro Technic Product, BD-20V) could be positioned close to the petri dish surface.
-The plasma treatment modifies the surface energy of the petri dish by breaking chemical bonds on the surface and introducing polar functional groups, thereby inducing the hydrophilic state [26].
+Plasma treatment was applied for 30 seconds in a simple custom-device with a closed chamber in PMMA and under ambient conditions.
+Plasma was generated with a BD-20AV high-frequency generator (230 V, ETP, IL, USA) with an electrode positioned close to the petri dish surface.
+The plasma treatment modifies the surface energy of the petri dish by breaking chemical bonds on the surface and introducing polar functional groups, thereby inducing the hydrophilic state [Al-Azzam2022].
 Approximately 3 µL of the FND suspension was then carefully deposited onto the clean petri dish using a micropipette.
@@
 After the solution had completely evaporated, the sample was ready for the optical experiments described in the following.
-The preparation of FNDs in the liquid environment also starts with the dilution steps and cleaning of the petri dish, but to maintain the FNDs in solution, the petri dish was not plasma treated and the solution was not left for evaporation.
-FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments [27].
+The preparation of FNDs for experiments in the liquid environment also starts with the dilution steps and cleaning of the petri dish, but to maintain the FNDs in solution, the petri dish was not plasma treated and the solution was not left for evaporation.
+FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments [Niora2024].
 The FNDs (Sigma-Aldrich 1003626896) were prepared following the same protocol.
 A suspension of FNDs was created by diluting 1 µL of the commercial stock solution in 1000 µL of deionized water, followed by sonication for 10 minutes to ensure a homogeneous dispersion.
-Approximately 500 µL of the suspension was deposited in a Petri dish that was thoroughly cleaned.
+Approximately 50 µL of the suspension was deposited in a sterile Petri dish that was freshly unwrapped.
 II.2 Optical Setup
-The dried FND sample was placed inside a mini-incubator (pecon XS 2000) in an optical setup based on a customized Thorlabs OTKB/M.
-Customization included a different objective (Nikon, CFI PLAN APO LAMBDA 100x OIL NA 1.45 WD 0.13) and long-distance condenser lens (Mitutoyo M Plan Apo 10x / 0.26).
-The setup is constructed around an inverted microscope configuration, equipped with a high-resolution camera (Teledyne Prime 95B), and two lasers for the simultaneous illumination and manipulation of the FNDs.
-A green laser (Coherent Verdi G5) with a wavelength of 532 nm was utilized for the excitation of the NV centers, enabling the initialization and readout of their spin states.
-The second laser (Cobolt Rumba, 3W) is operating in the near-infrared (NIR) at a wavelength of 1064 nm and used for FND trapping.
+The dried FND sample was placed inside a mini-incubator (pecon XS 2000) in an optical setup based on a customized Thorlabs OTKB/M modular optical tweezer system.
+Customization included a different objective and long-distance condenser lens.
+Vendors and other details are listed in Table 2 in the Appendix.
+The setup is constructed around an inverted microscope configuration, equipped with a high-resolution camera that capture wide-field bright-field images of the FND, and two lasers for the simultaneous illumination and manipulation of the FNDs.
+A green laser with a wavelength of 532 nm was utilized for the excitation of the NV centers, enabling the initialization and readout of their spin states.
+The second laser is operating in the near-infrared (NIR) at a wavelength of 1064 nm and used for FND trapping.
 This wavelength falls within the biological transparency window, which minimizes absorption and reduces the risk of significant heating effects in biological samples.
-The two lasers were aligned and their beams combined using a dichroic mirror, ensuring they shared the same optical path before being focused onto the sample through the microscope objective, as detailed in Fig.
-2.
-By trapping an FND and then adjusting the position of the green laser to maximize the signal on the photon counter, we ensured the same focal point for both the green and NIR lasers.
-Additionally, acousto-optic modulators (AOMs) were used to precisely control the timing and intensity of the laser beams, enabling fine-tuned manipulation of the FNDs.
+The two lasers were aligned and their beams combined using a dichroic mirror, ensuring that they shared the same optical path before being focused onto the sample through the microscope objective, as detailed in Fig.
+2a).
+By trapping an FND and then finetuning the position of the green laser to maximize the signal on the photon counter, we ensured the same focal point for both the green and NIR lasers.
+Additionally, acousto-optic deflectors were used to precisely control the timing and intensity of the laser beams by using Qudi software, enabling fine-tuned manipulation of the FNDs.
 As illustrated in Fig.
-2, several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection.
-A motorized flip-mirror ensures the ability to rapidly shift between camera detection or either spectrometer or photon count detection.
-The fluorescence spectrum of the FNDs is recorded and analyzed by a fiber coupled spectrometer (Ocean Optics X00428) while alternatively photon counts are measured with a fiber coupled avalanche photo detector (Excelitas SPCM-AQRH-13-Fc).
-In addition, to measure the ODMR spectrum of FNDs, an Agilent N5181A signal generator connected to an amplifier (MiniCircuits ZHL-45) is connected to a copper wire antenna placed approximately 1mm above the sample.
+2a), several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection.
+A motorized flip-mirror ensures the ability to rapidly shift between camera detection and either spectrometer or photon count detection.
+The fluorescence spectrum of the FNDs is recorded and analyzed by a fiber coupled spectrometer while alternatively photon counts are measured with a fiber coupled avalanche photo detector.
+In addition, to measure the CW ODMR spectrum of FNDs, a signal generator connected to an amplifier is connected to a single-turn copper loop antenna (5 mm) placed approximately 1 mm above the sample.
 When the flip-mirror is positioned towards the camera, the latter captures high-resolution images of the FNDs illuminated during the experiments in real time.
 The combination of different detection methods thus made it possible to continuously monitor the position of the FNDs, determine whether they were isolated or in aggregates, and analyse their photoluminescence properties throughout the experimental process.
-Figure 2: Illustration of the optical setup with the most important lasers, optical lenses, filters and equipment.
+Figure 2: a) Illustration of the optical setup with the most important lasers, optical lenses, filters and equipment.
+Details of the components are specified in Table 2 in the appendix.
+b) Laser sequence for fluorescence relaxometry measurement.
+In green the 532 nm laser and in red the 1064 nm laser.
+Pulse-widths are $t_{\rm trap}=20\mu s$, $t_{ini}=35\mu s$, $t_{RO}=15\mu s$ while the variable waiting time $\tau$ allows to record the relaxometry at times spanning from $7.2\mu s$ to $720\mu s$.
 II.3 FND Trapping
-For the FND trapping experiments, solution samples were used to ensure the FNDs remained mobile during the experiment.
-The NIR laser was utilized for trapping the FNDs in the focused beam [28, 29, 30, 31].
-For $T_{1}$ relaxometry, the temporal sequence of the green and trapping lasers, as illustrated in Fig.
-3, is carefully designed to minimize the overlap between the two lasers [19].
-This approach reduces potential spin-mixing effects from the NIR laser, while maintaining the stability of the trap.
-The green laser was pulsed for 35 µs for the excitation and 15 µs for the readout of the NV center spin states.
-The NIR laser was pulsed for 20 µs for the trapping of the FNDs.
-By using this sequence, we optimized the trapping stability during $T_{1}$ relaxometry measurement and minimized NIR irradiation potential impact on the measurements.
-The sensing experiments utilized the FND trapping protocol described below, and the CW NIR laser power was set to 60 mW to stabilize the FNDs in the optical trap.
-However, for manipulations involving changes to the surrounding medium, the CW NIR laser power was increased to 100 mW to ensure that the FND remained trapped despite turbulence in the medium.
-After the modifications, the laser power was reduced back to 60 mW for subsequent measurements.
-Figure 3: Laser sequence for T1 measurement. In green the 532 nm laser and in red the 1064 nm laser. $t_{NIR}=20\mu s$, $t_{I}=35\mu s$, $t_{RO}=15\mu s$ and $\tau$ from $7.2\mu s$ to $720\mu s$.
+All trapping experiments were performed in a fluid chamber containing 50 µL of the FND aqueous suspension unless stated otherwise.
+The NIR laser was utilized for trapping the FNDs in the focused beam [Neuman2004, Gennerich2017, Jones2015].
+For relaxometry, the temporal sequence of the green and trapping lasers, as illustrated in Fig.
+2b), avoids temporal overlap between the two lasers while maintaining the FND stably trapped [Russell2018a].
+Strictly speaking, this protocol does not constitute a conventional spin-relaxometry measurement.
+With an all-optical protocol, the method inherently involves potential inaccuracies due to photoionization induced by the green laser itself, thus affecting the NV- charge states.
+To address this limitation, measurement protocols incorporating microwave pulses explicitly designed to disentangle spin dynamics from photoionization effects have been developed.
+Examples include the microwave $\pi$-pulse method introduced by Ref.
+[Jarmola2023], as well as alternative sequences described in Ref.
+[Cardoso2023].
+We hence will refer to the measured decay time as $T_{\rm relax}$ throughout the remainder of this paper.
+The green laser was pulsed with initialization time $t_{ini}$ of 35 µs, and read-out time $t_{RO}$ of 15 µs, for the readout of the NV center fluorescence, while the NIR laser was on to trap the FNDs in sequences with duration $t_{\rm trap}$ of 20 µs.
 II.4 Sensing Experiments
-II.4.1 Gd3+ Magnetic Noise Detection
 For the experiments detecting paramagnetic species, a gadolinium solution (0.45 M in acetate) was gradually added to the FND’s environment.
 The concentrations were incrementally adjusted, and measurements were performed after each addition to evaluate the changes in the NV center properties.
-II.4.2 pH Sensing
 For the pH sensing experiments, an initial 0.1 M NaOH(aq) solution at pH 13 was prepared.
@@
 After each pH adjustment, the properties of the NV centers were measured while the FNDs were maintained in the optical trap at 60 mW.
-II.4.3 Temperature Sensing
 For the temperature sensing experiments, the medium containing the FNDs was initially maintained at 37 C using an incubator.
@@
 After each change of the set temperature of the incubator, we waited 10 minutes before the measurements were conducted.
+For the sensing of paramagnetic species and pH, changes to the surrounding medium was involved.
+To ensure that the FND remained trapped despite disturbances in the medium, the CW NIR laser power was temporarily increased to 100 mW while liquid was added.
+Subsequently, the laser power was reduced back to 60 mW for measurements with the new conditions, and the same FND.
 III Results
-We begin by presenting the main results of our experiments, starting with the examination of the general fluorescence properties of NV-centers within the FNDs through Photoluminescence (PL) spectroscopy.
+We begin by presenting the main results of our characterisation experiments, starting with the examination of the general fluorescence properties of NV-centers within the FNDs through photoluminescence (PL) spectroscopy.
 This foundational assessment provides an overview of the baseline optical characteristics and helps identify any immediate changes induced by the NIR laser irradiation.
-Following the PL spectrum analysis, we delve into the variations in fluorescence intensity as a function of optically detected magnetic resonance (ODMR) conditions.
+Following the PL spectrum analysis, we present the variations in fluorescence intensity as a function of optically detected magnetic resonance (ODMR) conditions.
 This step is crucial for understanding the impact of NIR irradiation on the magnetic resonance properties of the NV centers.
-Finally, we conduct T1-relaxometry to investigate the spin relaxation dynamics, providing a detailed examination of the spin-state populations and their evolution under different NIR laser parameters.
-This approach from general fluorescence to specific spin dynamics allows for a comprehensive evaluation of the NIR irradiation effects on the NV centers.
+Finally, we conduct fluorescence relaxometry to investigate the spin relaxation dynamics under different NIR laser parameters.
 PL spectrum measurement and ODMR measurement were conducted with CW NIR laser irradiation.
+The experiments conducted under dry conditions primarily serve as a thermal control benchmark.
+This dry setup facilitates the understanding of thermal phenomena, such as localized heating effects, by providing a clear reference to compare with conditions involving FNDs suspended in aqueous environments.
+We acknowledge that the dry condition itself lacks direct practical application in sensing contexts.
+For each set of experiments described in the subsections below, we exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW.
 III.1 Photoluminescence Spectrum Analysis
-We conducted PL spectroscopy of FNDs with varying NIR power. The FNDs were either dried on glass substrates or trapped in water to examine the influence of the surrounding medium.
-Effect of NIR Laser Power on PL Spectrum Intensity
-We exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW and for each power observe the PL spectrum from 550 nm to 750 nm with our spectrometer.
+As the NV neutral and negative charge states have distinct optical emission spectra, we conducted PL spectroscopy of FNDs at varying NIR power to investigate its impact.
+The FNDs were either dried on glass substrates or trapped in water to examine the influence of the surrounding medium.
+We record the PL spectrum in the range between 550 nm to 745 nm with the spectrometer for each of the different values of the CW NIR laser power.
 Figure 3: Effect of CW NIR laser on dry FND emission spectrum.
-Optical emission spectrum of dry FND on glass under CW NIR laser exposure at 0 mW (green), 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
-Our PL measurements revealed significant alterations in the spectral intensities of both zero-phonon lines (ZPL) from the two NV charge states upon CW NIR laser exposure.
+Optical emission spectrum of one FND, dry on glass, under CW NIR laser exposure at 0 mW (green), 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
+Upon CW NIR laser light exposure, our PL measurements revealed a significant drop in the total PL intensity.
 Notably, as a function of NIR laser power there was a slightly more pronounced reduction in the spectral intensity of NV- compared to NV0.
-As shown in Fig.
-4, the ZPL intensity of NV0 at 575 nm decreased by 29.5% and 32.4%, while the NV- ZPL intensity at 637 nm decreased by 52.8% and 62.3% when the CW NIR laser power was increased from 0 mW to 30 mW and 60 mW, for the condition with FNDs dry on glass.
 For the fluorescence spectrum (like Fig.
-4), the curves remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30mW to 60mW (data not shown) and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water, cf.
-Fig.
-5.
-Quantitatively, we observe a 1.2% general intensity difference from 30 mW to 60 mW for FNDs on glass with water while the same figure is 0.7% for FNDs trapped in water.
-Figure 5: Effect of CW NIR laser on trapped FND emission spectrum. Optical emission spectrum of FND trapped in water CW NIR laser at 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
-To determine the ratio of NV0 to NV- emission intensities under CW NIR laser powers of 0 mW, 30 mW, and 60 mW we integrated the emission intensities over the respective spectral ranges for NV0 (550–620 nm) and NV- (655–750 nm), but compared to the procedure of Ref.
-[13], we applied a normalization factor to account for the difference in spectral ranges:
-$\frac{\text{NV}^{0}}{\text{NV}^{-}}=\left(\frac{\sum_{\lambda=550}^{620}I_{\lambda}}{\sum_{\lambda=655}^{750}I_{\lambda}}\right)\times\frac{95}{70}$ (1)
-We observe that the NV0/NV- emission ratio stays the same for trapped FNDs, 0.45 for both 30mW and 60mW NIR laser power.
-In contrast, it rises from 0.39 without NIR irradiation to 0.51 with 60 mW NIR laser power for dry FND on glass.
+3), the curves remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30mW to 60mW (data not shown) and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water.
+To quantify the changes in relative charge state emission intensities under CW NIR laser irradiation, we determine the ratio $\sigma_{0-}$ between NV0 and NV- emission intensities by integrating over the spectral ranges for NV0 (550–620 nm) and NV- (655–745 nm), respectively.
+Compared to the analysis used in Ref.
+[Sow2020], we applied a normalization factor 19/14 to account for the difference in the relative spectral ranges:
+$\sigma_{0-}\equiv\left(\frac{\sum_{\lambda=550}^{620}I_{\lambda}}{\sum_{\lambda=655}^{745}I_{\lambda}}\right)\times\frac{19}{14}$ (1)
+For FNDs trapped with NIR powers of 30 mW and 60 mW, we observe an averaged similar emission ratios $\sigma_{0-}$ of $0.48\pm 0.09$.
+In contrast, for FNDs dried on glass, we measure an emission ratio of $0.43\pm 0.11$ without NIR irradiation that increases to $0.57\pm 0.15$ for a NIR laser power of 60 mW NIR.
 All results are displayed in Fig.
-10.
+6 parts a) and c).
 III.2 Optically Detected Magnetic Resonance Measurements
-We further examined the influence of CW NIR laser power on the ODMR resonance frequency and contrast for FNDs dried on glass substrates and those trapped in water.
-Effect of CW NIR Laser Power on the Zero Field Splitting of the NV Center
-We exposed the FNDs to CW NIR laser powers of 0 mW, 30 mW, and 60 mW, and measured the resulting ODMR spectra.
-Our primary observation was a shift in the resonance frequency with increasing CW NIR laser power, as illustrated with an example in Fig.
-6.
+In the next step, we elaborate on the influence of CW NIR laser light on the ODMR resonance frequency and contrast, revealing the impact of the NIR trapping laser on the NV electron spin properties.
+The ODMR spectra were acquired in continuous-wave (CW) mode with constant intensities of both the 532 nm pump and 1064 nm trapping lasers and the microwave source, while sweeping the frequency of the microwave in steps of 0.3 MHz.
+Generally and for all FNDs, we observe a redshift in the resonance frequency with increasing CW NIR laser power, as illustrated in Fig.
+4 by the spectra from one FND with 0 mW and 60 mW NIR power.
 Figure 4: Effect of CW NIR laser power on ODMR spectrum of dry FND.
-ODMR spectra of dry FNDs under CW NIR laser exposure at 0 mW (green) and 60 mW (red) of 1064 nm laser power.
+ODMR spectra of one FND under CW NIR laser exposure at 0 mW (green) and 60 mW (red) of 1064 nm laser power.
 The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed.
+The power of the 532 nm laser was 0.7 mW, measured before the objective lens, and the microwave amplifier was set to -2dB.
 For FNDs dried on glass substrates, the resonance frequency shifted significantly to the lower frequencies.
-Specifically, the resonance frequency shift increased in absolute value from $-0.31\pm 0.54$ MHz at 0 mW to $-5.83\pm 046$ MHz at 60 mW of CW NIR power; both values given relative to the expected 2.87 GHz.
-In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $-0.84\pm 0.21$ MHz at 30 mW and $-1.58\pm 0.33$ MHz at 60 mW (relative to the expected 2.87 GHz).
+Taking 2.87 GHz as reference value for the zero field splitting, the resonance frequency shift increased in absolute value from $0.31\pm 0.54$ MHz at 0 mW to $5.83\pm 046$ MHz at 60 mW of CW NIR power.
+In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $0.84\pm 0.21$ MHz at 30 mW and $1.58\pm 0.33$ MHz at 60 mW.
 Examples are shown in Fig.
-6 and Fig.
-7 while combined results are displayed in Fig.
-11.
+4 and combined results displayed in Fig.
+7.
 As before, for the similar experiment when water was added to the same FND that remained stuck on the glass slide, the ODMR spectra show a trend very similar to Fig.
 4 although the difference between 30 mW and 60 mW of NIR power is less pronounced (data not shown).
-Figure 7: Effect of CW NIR laser power on ODMR spectrum of trapped FND.
-ODMR spectra of FNDs trapped in water under CW NIR laser exposure at 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
-The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed.
-III.3 T1 Relaxometry Measurements
-We conducted $T_{1}$ relaxometry experiments under different NIR laser power.
-As above, FNDs were either dried on glass substrates or suspended in water to assess the influence of the surrounding medium on the T1 relaxometry measurements with NIR irradiation.
-The samples were exposed to NIR laser powers of 0 mW, 30 mW, and 60 mW, and the corresponding $T_{1}$ values were measured.
-For these experiments, the pulsing sequence in Fig.
-3 was applied.
-Effect of NIR Laser Power on $T_{1}$ Relaxation Times
-First, we investigate how the T1 value of dry FNDs is affected by NIR irradiation.
-For FNDs dried on glass substrates, we observed a significant decrease in the $T_{1}$ relaxation times with increasing NIR laser power.
-As shown in the example in Fig 8, the mean $T_{1}$ value decreased from $241\pm 21~{}\mu$s at 0 mW NIR power to $136\pm 19~{}\mu$s at 60 mW NIR power.
-When water was added to the same FND that remained stuck on the glass slide, the $T_{1}$ values showed very similar changes.
-The mean $T_{1}$ value decreased from $237\pm 25~{}\mu$s at 0 mW NIR power to $154\pm 17~{}\mu$s at 30 mW NIR power and $143\pm 21~{}\mu$s at 60 mW NIR power.
-(data not shown).
-Figure 8: Effect of NIR laser power on $T_{1}$ relaxation times of dry FNDs.
-$T_{1}$ relaxometry measurements for FNDs dried on glass under NIR laser powers of 0 mW (green) and 60 mW (red) at 1064 nm wavelength.
-Pulsing as in Fig 3.
-For FNDs trapped in water, the decrease in $T_{1}$ due to NIR irradiation was significantly less.
-As depicted in the example in Fig.
-9, we do not observe any significant change in the 2 curves.
-The $T_{1}$ values for trapped FNDs were $179\pm 8~{}\mu$s at 30 mW and $176\pm 8~{}\mu$s at 60 mW, showing a negligible decrease compared to the values at 0 mW, with all results summarized in Fig.
-11.
-Figure 9: Effect of NIR laser power on $T_{1}$ relaxation times of trapped FNDs.
-$T_{1}$ relaxometry measurements for FNDs trapped in water under pulsed NIR laser powers of 30 mW (orange) and 60 mW (red) at 1064 nm wavelength.
-III.4 Combined Statistical Analysis
-In this section, we present a comprehensive statistical analysis of our experimental results.
-Relationship Between NV0/NV- Ratio and ODMR Frequency Shift
-Figures 10.a, 10.c, 11.a and 11.b illustrate the dependence of the NV0/NV- fluorescence intensity ratio and the ODMR resonance frequency shift for FNDs under varying CW NIR laser powers.
-For FNDs dried on glass substrates, we observe a significant increase in the NV0/NV- fluorescence intensity ratio and a corresponding shift in the ODMR resonance frequency with increasing CW NIR power from 0 mW to 60 mW.
-In contrast, for FNDs trapped in water, the NV0/NV- fluorescence intensity ratio and ODMR frequency shift exhibit much smaller changes with increasing CW NIR power.
-This suggests that the aqueous environment mitigates the effects of CW NIR irradiation.
-Correlation Between $T_{1}$ Relaxation Times and ODMR Frequency Shift
-Figures 10.b, 10.d, 11.a and 11.c show the relationship between the $T_{1}$ relaxation times and the ODMR resonance frequency shift for FNDs under varying NIR laser powers.
-$T_{1}$ was obtained with the pulse sequence as in Fig 3 while the ODMR was obtained with CW NIR irradiation.
-For FNDs dried on glass substrates, there is an effect between $T_{1}$ times and ODMR frequency shifts with increasing NIR power.
-The significant decrease in $T_{1}$ times and increase in frequency shift suggest that NIR irradiation enhances spin relaxation processes.
-For FNDs trapped in water, the $T_{1}$ relaxation times remain relatively constant, and the ODMR frequency shifts exhibit minimal changes with increasing NIR power.
-Overall, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts.
+[…]
+III.3 Relaxometry Measurements
+Last, we conducted relaxometry experiments under different NIR laser power, to extract values for the corresponding $T_{\rm relax}$.
+As above, FNDs were either dried on glass substrates or suspended in water to assess the influence of the surrounding medium on the relaxometry measurements with NIR irradiation.
+For these experiments, the all-optical pulsing sequence in Fig.
+2b) was applied.
+[…]
+Figure 5: Effect of NIR laser power on $T_{\rm relax}$ relaxation times of dry FNDs.
+$T_{\rm relax}$ relaxometry measurements for one FND dried on glass under NIR laser powers of 0 mW (green) and 60 mW (red) at 1064 nm wavelength.
+Pulsing as in Fig 2b).
+[…]
+Overall, as showed in Fig.
+6 and Fig.
+7, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts.
 Further, our results confirm that with specific NIR irradiation sequences and optimizing laser parameters, cf.
 Fig.
-3 and Ref.
-[19], potential detrimental effects of NIR on $T_{1}$ relaxation can be significantly mitigated, even with FNDs dry on glass.
-Figure 10: a) Correlation between NIR laser power, ODMR resonance frequency shift and NV0/NV- fluorescence intensity ratio for FNDs trapped inside water.
-b) Correlation between NIR laser power, ODMR resonance frequency shift and T1-relaxometry times for FNDs trapped inside water.
-c) Correlation between NIR laser power, ODMR resonance frequency shift and NV0/NV- fluorescence intensity ratio for FNDs dry on glass.
-d) Correlation between NIR laser power, ODMR resonance frequency shift and T1-relaxometry times for FNDs dry on glass.
+2b) and Ref.
+[Russell2018a], potential detrimental effects of NIR on $T_{\rm relax}$ relaxation can be significantly mitigated.
+Figure 6: a) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs trapped inside water.
+b) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs trapped inside water.
+c) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs dry on glass.
+d) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs dry on glass.
 Data for 30 FNDs exposed to 0 mW (green), 30 mW (orange), and 60 mW (red) of 1064 nm NIR laser irradiation.
-CW NIR for ODMR and spectra, pulsed NIR as in Fig 3 for T1-relaxometry.
-In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and no such trend for FNDs in the optical trap in water.
+CW NIR for ODMR and spectra, pulsed NIR as in Fig 2b) for fluorescence relaxometry.
+In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and and no or much less pronounced trend for FNDs in the optical trap in water.
 Figure 7: Summary of experimental results, showing variation for individual nanodiamonds.
-(a) ODMR resonance frequency shift, (b) NV0/NV- fluorescence intensity ratio, and (c) $T_{1}$ relaxometry times as functions of NIR laser power.
+(a) ODMR resonance frequency shift, relative to the expected resonance at 2.87 GHz, (b) fluorescence intensity ratio $\sigma_{0-}$, and (c) $T_{\rm relax}$ relaxometry times as functions of NIR laser power.
 Data for 5 random FNDs dry on glass (black stars) and trapped in water (gray circles) with 1064 nm NIR laser irradiation.
-Statistical Analysis Using ANOVA
-To study the significance of the observed effects, we performed an Analysis of Variance (ANOVA) test on the $T_{1}$ relaxation times, ODMR frequency shifts, and NV0/NV- fluorescence intensity ratios.
-The ANOVA results are summarized in Table 1.
+To study the significance of the observed effects, we performed an Analysis of Variance (ANOVA) test on the $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
+The ANOVA results are summarized in Table 1 (one-way Kruskal–Wallis test, $\alpha=0.05$).
 Table 1: ANOVA test results for the effect of NIR laser power on FNDs dried on glass substrates, on glass after adding water and trapped in water.
-The p-values indicate the significance of differences observed across NIR power levels for $T_{1}$ relaxation times, ODMR frequency shifts, and NV0/NV- ratios.
+The p-values indicate the significance of differences observed across NIR power levels for $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
 p-value
 FNDs dried on glass
-$T_{1}$ Relaxation Time $<0.0001$
-ODMR Frequency Shift $<0.0001$
-NV0/NV- Ratio 0.0024
+$T_{\rm relax}$ Relaxation Time <0.0001<0.0001
+ODMR Frequency Shift <0.0001<0.0001
+$\sigma_{0-}$ Ratio 0.0024
 FNDs on glass with water
-$T_{1}$ Relaxation Time $0.072$
+$T_{\rm relax}$ Relaxation Time $0.072$
 ODMR Frequency Shift $0.008$
-NV0/NV- Ratio 0.201
+$\sigma_{0-}$ Ratio 0.201
 FNDs Trapped in Water
-$T_{1}$ Relaxation Time 0.221
+$T_{\rm relax}$ Relaxation Time 0.221
 ODMR Frequency Shift 0.073
-NV0/NV- Ratio 0.371
-For FNDs dried on glass substrates, the very low p-values (p $<0.01$) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant.
+$\sigma_{0-}$ Ratio 0.371
+For FNDs dried on glass substrates, the very low p-values (p <0.01<0.01) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant.
 This indicates that NIR irradiation has a considerable impact on the NV centers’ properties in air.
-For FNDs on glass substrates after adding water, the medium p-values ($0.20.01<p<$ 0.2) for all three parameters confirm that the differences observed with varying NIR laser powers are not always statistically significant.
+For FNDs on glass substrates after adding water, the mean p-values ($0.01<p<0.2$) for all three parameters confirm that the differences observed with varying NIR laser powers are not always statistically significant.
 In fact, the difference between no NIR laser and NIR laser is significant but the differences between 30 mW and 60 mW power of NIR laser are not significant.
-In contrast, for FNDs trapped in water, the p-values are high (p $>0.07$), suggesting that the differences observed across different NIR power levels are not statistically significant.
+In contrast, for FNDs trapped in water, the p-values are high (p >0.07>0.07), suggesting that the differences observed across different NIR power levels are not statistically significant.
 This reinforces the conclusion that the aqueous environment mitigates the effects of NIR irradiation on the NV centers.
-III.5 Charge State Dynamics
-To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model [32].
+III.4 Model of Charge State Dynamics
+To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model [Qian2022], supplemented with insights from [Meirzada2018].
 This model accounts for the various optical transitions, inter-system crossings, ionization, and recombination processes that occur when NV centers are simultaneously exposed to green and NIR laser excitation.
 The NV center can exist in two charge states: the negatively charged NV- and the neutrally charged NV0.
-Each charge state has its own ground and excited electronic states, as well as inter-system crossing pathways to metastable singlet states.
+Each charge state has its own ground and excited electronic states, as well as inter-system crossing pathways to metastable quartet states.
 The model considers 8 states as illustrated in the schematic energy level diagram in Fig.
-12.
-The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state [32].
-The rates of change had to be adapted to our experimentation with green 532 nm and NIR 1064 nm laser.
+8.
+The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state [Qian2022].
+The rates of change were adapted to our experimental situation with just two lasers, the green 532 nm and the NIR 1064 nm laser, following [Meirzada2018].
+Values used in this work for the rates are provided in the appendix, Table 3.
 The model is based on the following assumptions:
@@
   • No External Fields: Effects of external magnetic or electric fields on the transition rates are not included.
-Figure 12: Energy level diagram of the NV center, showing the ground states (P1, P2 and P7), excited states (P3, P4 and P8), singlet states (P5 and P6), and the transitions between them.
+Figure 8: Energy level diagram of the NV center, showing the ground states (P1, P2 and P7), excited states (P3, P4 and P8), quartet states (P5 and P6), and the transitions between them.
 Solid arrows represent radiative transitions, dotted arrows represent non-radiative transitions, and dashed arrows represent ionization and recombination processes.
-Population Dynamics of NV- and NV0 Centers
 The simulation results reveal that at low CW NIR laser powers, the population of NV- centers increases slightly, while the NV0 population decreases.
@@
 This inversion is due to the saturation of recombination processes and the dominance of ionization processes at higher CW NIR powers, which ionize NV- centers back to NV0.
-The threshold depends on the recombination rate, ionization rate and on the initial population of each state.
-Fluorescence Intensity and NV0/NV- Ratio
 The total fluorescence intensity from the NV centers decreases with increasing CW NIR power (Fig.
-13).
+9).
 This decrease is primarily due to the reduced population of NV- centers and the lower fluorescence quantum yield of NV0 centers.
 The fluorescence from NV- centers dominates the emission under low CW NIR power, but as the CW NIR power increases, the contribution from NV0 centers becomes more significant and we observe a saturation in the fluorescence.
-The ratio of NV0 to NV- fluorescence intensities ($F_{\text{NV}^{0}}/F_{\text{NV}^{-}}$) increases with CW NIR power.
+The $\sigma_{0-}$ ratio increases with CW NIR power.
 At low CW NIR powers, this ratio is low due to the predominance of NV- fluorescence.
@@
 This change in the fluorescence ratio affects the overall spectral characteristics and can influence the detection sensitivity in applications relying on NV- fluorescence.
-Consistency with Experimental Observations
 The trends observed in the simulation align with the experimental results.
-The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity and NV0/NV- fluorescence intensity ratio observed experimentally.
-The simulation provides a quantitative framework to understand how NIR illumination influences the charge state dynamics and, consequently, the optical and spin properties of NV centers.
-Figure 13: Total fluorescence intensity from NV centers as a function of NIR laser power. (Black: Theory, Red: Experimental observations.
-III.6 Temperature Simulation
+The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity (Fig.
+9) and $\sigma_{0-}$ ratio (Fig.
+6a,c)) observed experimentally.
+Figure 9: Total fluorescence intensity from NV centers as a function of NIR laser power.
+(Black: Theory, Red: Experimental observations with errorbars, obtained as a statistical average over results from 10 different FNDs with $\sigma_{0-}$ within the range of $\sigma_{0-}=0.42\pm 0.04$.)
+III.5 Temperature Simulation
 To model the temperature increase induced by the NIR laser on FNDs, we simulated the heat transfer.
 The simulation accounts for the absorption of laser energy by the FND and the subsequent heat dissipation into the surrounding medium (water or air).
-The key processes considered are:
-  1. 1.
-    Energy Absorption: The FND absorbs a portion of the incident laser power, leading to an increase in its internal energy and temperature.
-  2. 2.
-    Heat Loss: Heat is lost from the FND to the surrounding medium through conduction and convection, depending on the thermal properties of the medium.
+To quantify the temperature rise produced by the 1064 nm laser on a spherical fluorescent nanodiamond (radius $R=35\;\text{nm}$), we solve Fourier’s heat‑conduction equation with an internal absorption source inside the particle and pure conduction in the surrounding infinite medium (water or air):
+$\rho_{d}c_{d}\frac{\partial T_{d}}{\partial t}$ $=\frac{1}{r^{2}}\frac{\partial}{\partial r}\!\left(k_{d}r^{2}\frac{\partial T_{d}}{\partial r}\right)+q_{\mathrm{abs}}(r),$ $0\leq r<R,$
+$\rho_{m}c_{m}\frac{\partial T_{m}}{\partial t}$ $=\frac{1}{r^{2}}\frac{\partial}{\partial r}\!\left(k_{m}r^{2}\frac{\partial T_{m}}{\partial r}\right),$ $r>R,$
+subject to the interface and far‑field conditions
+$T_{d}(R,t)=T_{m}(R,t),$
+$k_{d}\!\left.\frac{\partial T_{d}}{\partial r}\right|_{r=R}=k_{m}\!\left.\frac{\partial T_{m}}{\partial r}\right|_{r=R},\qquad\lim_{r\to\infty}T_{m}(r,t)=T_{\infty}.$
+Here $\rho$, $c$, and $k$ are the density, specific heat, and thermal conductivity of the nanodiamond and the surrounding medium.
+The laser absorption is treated as a uniform volumetric heat source inside the FND,
+$q_{\mathrm{abs}}(r)=\begin{cases}\dfrac{P_{\mathrm{abs}}}{\tfrac{4}{3}\pi R^{3}},&r<R,\\[6.0pt] 0,&r\geq R,\end{cases}$
+where $P_{\mathrm{abs}}=\sigma_{\mathrm{abs}}I_{0}$ is the absorbed optical power for a beam of intensity $I_{0}$ and absorption cross‑section $\sigma_{\mathrm{abs}}$.
 The simulation is based on the following assumptions:
@@
   • Constant Thermal Properties: The thermal properties of the FND and the medium are assumed constant over the temperature range studied.
-  • Steady-State Heat Transfer Coefficient: The convective heat transfer coefficient $h$ is considered constant and based on the thermal conductivity of the medium.
+  • Steady-State Heat Transfer Coefficient: The convective heat transfer coefficient is considered constant and based on the thermal conductivity of the medium.
   • No Phase Change: The medium remains in the same phase (liquid for water, gas for air) throughout the simulation.
-Temperature Evolution During a 500 ns NIR Pulse
 […]
@@
   • The temperature increases during the laser pulse and decreases more slowly than in water after the laser is turned off.
-Consistency with Experimental Observations
-The simulation results align with experimental observations where pulsed NIR illumination causes negligible thermal effects in aqueous environments.
-This supports the conclusion that charge state dynamics are the primary mechanism affecting NV center behavior under short NIR pulses in water.
+The simulation results align with experimental observations where pulsed NIR illumination causes negligible thermal effects in aqueous environments, cf.
+Fig.
+7.
 In experiments conducted in air, temperature increases contribute to shifts in ODMR frequency and should be accounted for.
 These shifts correspond to a 2 MHz shift in the ODMR resonance frequency which is what we experimentally observe.
-IV Discussion: Effects of NIR Light on FNDs
-Our experimental and simulation results reveal a complex interplay between thermal dynamics and charge state transitions in NV within FNDs under NIR illumination.
-The observed effects include a decrease in total fluorescence intensity, reduction in spin-lattice relaxation time, shift in the ODMR resonance frequency, decrease in ODMR contrast, and an increase in the NV0/NV- fluorescence ratio.
-IV.1 Thermal Effects Induced by NIR Illumination
-One primary mechanism contributing to these observations is the localized heating induced by the NIR laser.
-The shift in the ODMR resonance frequency with increasing NIR power, as shown in Figures 6, 7 and 10, and 11a), suggests an increase in temperature within the FNDs.
-This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately $-74$ kHz/K [33, 34].
-However, the extent of the thermal effect is significantly influenced by the surrounding medium’s thermal properties.
-Media with higher thermal conductivity and heat capacity, such as water, dissipate heat more efficiently than air, resulting in smaller temperature increases in the FNDs.
-This is supported by our temperature simulations, which show minimal temperature rise in water compared to air during NIR illumination.
-Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts and negligible changes in $T_{1}$ times with increasing NIR power, as shown in Figures 7, 10 and 11.
-IV.2 Charge State Dynamics and Fluorescence Changes
-In addition to thermal effects, the NIR laser alters the charge state dynamics of the NV centers.
-Our simulations of charge state dynamics indicate that at low NIR power levels, there is a slight increase in the NV- population due to enhanced recombination processes facilitated by NIR photons.
-However, as the NIR power increases beyond a threshold (approximately 10 mW with our parameters), the NV- population begins to decrease while the NV0 population increases.
-This inversion is attributed to the saturation of recombination processes and increased ionization rates at higher NIR powers.
-These changes in charge state populations directly affect the fluorescence properties of the FNDs.
-The total fluorescence intensity decreases with increasing NIR power, as shown in Figure 13.
-This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers [35].
-The increase in the NV0/NV- fluorescence ratio with higher NIR power, depicted in Figures 10 and 11, further confirms the shift in charge state distribution.
-Additionally, the increased NV0 background fluorescence dilutes the ODMR contrast since NV0 centers do not exhibit spin-dependent fluorescence variations.
-This is reflected in the decreased ODMR contrast observed in our measurements (Figures 6-7).
-Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{1}$ times [20, 36].
-IV.3 Interplay Between Thermal Effects and Charge State Dynamics
-Although both thermal effects and charge state dynamics significantly impact the properties of the NV centers under NIR illumination, their relative contributions depend on the environmental conditions.
-In FNDs dried on glass substrates, the pronounced ODMR frequency shifts suggest that thermal effects dominate due to inefficient heat dissipation in air.
-In contrast, for FNDs trapped in water, the minimal temperature increase reduces thermal effects, making charge-state dynamics more prominent in influencing the NV centers’ behavior.
-The increase in the NV0/NV- ratio and the decrease of T1-relaxometry times with NIR power without significant ODMR frequency shifts (Figure 10) indicates that charge state transitions are the primary mechanism affecting the fluorescence and spin properties in aqueous environments.
-IV.4 Alternative Explanations and Considerations
-Although thermal effects and charge state dynamics provide a comprehensive explanation for our observations, alternative mechanisms should be considered.
-Multi-photon absorption processes could directly excite the NV centers without significant heating.
-However, the relationship between ODMR frequency shift and NIR power (Figures 7 and 10) supports the thermal interpretation in our experiments.
-Moreover, the significant differences in the effects of NIR irradiation between FNDs in air and water emphasize the critical role of the surrounding medium.
-The statistical analysis using ANOVA (Table 1) confirms that the differences observed in FNDs dried on glass are statistically significant, whereas those in FNDs trapped in water are not, reinforcing the influence of environmental thermal properties.
-V Sensing Capabilities of Trapped FNDs
+A direct comparison of Fig.
+6c)–d) with Fig.
+6a)–b) shows that the environment (air versus water) is the controlling parameter.
+In air, the ODMR resonance shifts by $5.8\pm 0.5$ MHz and $T_{\rm relax}$ falls by $(44\pm 6)\%$ when the CW NIR power is raised from 0 to 60 mW, whereas in water the corresponding changes are only $1.6\pm 0.3$ MHz and $(2\pm 5)\%$, respectively.
+By contrast, the NV0/NV- fluorescence intensity ratio $\sigma_{0-}$ responds similarly to NIR power in both media (Fig.
+6a,c; projection on the ordinate axis), confirming that charge-state kinetics are governed by photon flux rather than heat removal.
+For completeness we note that NV- centres possess a weak intersystem transition between the metastable singlet states $^{1}E\,(p_{6})\!\rightarrow\!{}^{1}A_{1}\,(p_{5})$ with a zero-phonon line at $\lambda_{\mathrm{ZPL}}=1042$ nm and a room-temperature homogeneous width of only a few GHz ($\lesssim 10^{-4}$ nm) [Acosta2010, Rogers2015].
+Our trapping beam at 1064 nm is therefore red-detuned by $\Delta\lambda=22$ nm, corresponding to $\Delta\nu\simeq 6$ THz ($\simeq$  25 meV), i.e.
+$\sim 10^{3}$ times the natural linewidth and well outside the vibrational sideband of the singlet manifold.
+Values for the measured absorption cross-sections decrease from $\sigma\!\approx\!2\times 10^{-19}$ cm2 on resonance to below $10^{-21}$ cm2 at this detuning [Acosta2010], so direct optical pumping of the singlet transition by the 1064 nm light is negligible.
+The effects reported in this work (PL quenching, ODMR shifts, $T_{\rm relax}$ variations) therefore arise from photothermal heating and photo-ionisation dynamics, not from resonant excitation of the 1042 nm line.
+IV Sensing Capabilities of Trapped FNDs
 As a final step in the comprehensive investigations, this section explores the potential of optically trapped FNDs with NV centers for sensing applications.
 Experiments focused on three key parameters: Gd3+ concentration, pH, and temperature.
-Measurements included ODMR, T1 relaxation, and optical spectrum analysis under variation of the three parameters mentioned above.
+Measurements included ODMR, fluorescence relaxation, and optical spectrum analysis under variation of the three parameters mentioned above.
+Results are summarized in Fig.
+10.
 Figure 10: Variation in three sensing parameters when subject to controlled variation of paramagnetic species, pH, and temperature.
@@
 b) ODMR resonance frequency shift VS pH value.
-c) ODMR resonance frequency shift VS Temperature increase.
-d) T1-relaxometry time VS Gd3+ concentration.
-e) T1-relaxometry time VS pH value.
-f) T1-relaxometry time VS Temperature increase.
-g) NV0/NV- fluorescence intensity ratio VS Gd3+ concentration.
-h) NV0/NV- fluorescence intensity ratio VS pH value.
-i) NV0/NV- fluorescence intensity ratio VS Temperature increase.
+c) ODMR resonance frequency shift (relative to the expected 2.87 GHz) VS Temperature increase (relative to 37 C).
+d) Fluorescence relaxometry time $T_{\rm relax}$ VS Gd3+ concentration.
+e) $T_{\rm relax}$ VS pH value.
+f) $T_{\rm relax}$ VS Temperature increase.
+g) $\sigma_{0-}$ fluorescence intensity ratio VS Gd3+ concentration.
+h) $\sigma_{0-}$ fluorescence intensity ratio VS pH value.
+i) $\sigma_{0-}$ fluorescence intensity ratio VS Temperature increase.
 The boxplot represent results for 10 FNDs that were trapped for each condition with 60 mW 1064 nm CW NIR laser.
-In each case, results for Gd3+ 0M, pH 13 and 0K trapped were used as reference values for the individual FNDs.
+In each case, results for 0 M Gd3+, pH 13 and 0 K trapped were used as reference values for the individual FNDs.
 The red squares emphasize the optimal measurement method for each sensing parameter.
-The ODMR resonance frequency value is given relative to the expected 2.87 GHz.
-V.1 pH Sensing
+IV.1 pH Sensing
 The impact of environmental pH on the properties of NV centers was evaluated using buffer solutions with pH values of 1, 7, and 13.
-Figure 15: NV0/NV- fluorescence intensity ratio during pH value changes.
-Measurement of the same FND while changing from pH13 to pH1 by adding HCl solution (Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
-  • T1 Relaxation: A decrease in T1 values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig.
-14.h).
-This is consistent with reports that high pH conditions stabilize spin relaxation dynamics [1].
-  • Optical Spectrum: A decrease in the NV0/NV- fluorescence intensity ratio with increasing pH (Fig.
-14.e and  15) suggests that alkaline conditions favor the NV- charge state, possibly due to deprotonation of surface functional groups, which enhances negative surface potential [37].
-The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser shows a clear advantage of using FNDs in suspension for sensing.
+Figure 11: $\sigma_{0-}$ fluorescence intensity ratio during pH value changes.
+Measurement of the same FND while changing from pH13 to pH1 by adding HCl solution.
+Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
+  • Fluorescence relaxation: A decrease in $T_{\rm relax}$ values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig.
+10.h).
+This is consistent with reports that high pH conditions stabilize spin relaxation dynamics [Fujisaku2019, Rendler2017].
+  • Optical Spectrum: A decrease in the $\sigma_{0-}$ ratio with increasing pH (Fig.
+10.e and Fig.
+11) suggests that alkaline conditions favor the NV- charge state.
+It may be attributed to the deprotonation of surface functional groups such as carboxyl (–COOH) and hydroxyl (–OH) groups present on the nanodiamond surface.
+Deprotonation of these groups under alkaline conditions generates negatively charged surface sites, thereby enhancing electron donation to nearby NV centers and favoring the stabilization of the NV- charge state [karaveli2016].
+The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (Fig.
+11) shows a clear advantage of using FNDs in suspension for sensing.
   • ODMR: No significant changes in resonance frequency (Fig. 10.b) were detected across the tested pH range, indicating minimal impact of pH on electronic spin states.
-The NV0/NV- fluorescence intensity ratio’s dependence on pH underscores the importance of surface chemistry in modulating charge state dynamics, especially with FNDs in suspension.
-V.2 Paramagnetic Species Sensing
+The $\sigma_{0-}$ ratio’s dependence on pH underscores the importance of surface chemistry in modulating charge state dynamics, especially with FNDs in suspension.
+To further explore the dependence of the NV charge state on pH, Fig.
+11 presents the evolution of the $\sigma_{0-}$ ratio when varying pH from 13 to 1 for the same FND, both under trapping conditions and without NIR laser exposure.
+The trapped FNDs in suspension exhibit a significantly stronger modulation of $\sigma_{0-}$ compared to the same particle measured on glass.
+This enhanced responsiveness highlights the improved surface accessibility and chemical exchange in solution, reinforcing the utility of optically trapped FNDs for pH nanosensing.
+The pronounced increase in NV center fraction under alkaline conditions (lower $\sigma_{0-}$) confirms the surface deprotonation hypothesis, while acidic conditions suppress this stabilization.
+Overall, these results validate that $\sigma_{0-}$ serves as a reliable and sensitive optical proxy for probing local pH variations in biological or environmental settings.
+IV.2 Paramagnetic Species Sensing
 The sensitivity of NV centers to paramagnetic gadolinium ions (Gd3+) was investigated through measurements at 3 different values of the concentration.
-Figure 16: NV0/NV- fluorescence intensity ratio during gadolinium concentration changes.
-Measurement of the same FND while changing gadolinium concentration from 0M to 0.1M by adding a gadolinium solution (0.45 M in acetate) (Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
-  • T1 Relaxation: A significant decrease in T1 values was observed with increasing Gd3+ concentration (Fig.
-14.g).
-This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment [38].
-  • Optical Spectrum: A decrease in the NV0/NV- fluorescence intensity ratio was observed (Fig.
-14.d and  16).
-This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 [39].
-The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser shows a clear advantage of using FNDs in suspension for sensing.
+Figure 12: $\sigma_{0-}$ fluorescence intensity ratio during gadolinium concentration changes.
+Measurement of the same FND while changing gadolinium concentration from 0M to 0.1M by adding a gadolinium solution (0.45 M in acetate).
+Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
+  • Fluorescence relaxation: A significant decrease in $T_{\rm relax}$ values was observed with increasing Gd3+ concentration (Fig.
+10.g).
+This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment [Martinez2020, Rendler2017].
+  • Optical Spectrum: A decrease in $\sigma_{0-}$ was observed (Fig.
+10.d and Fig.
+12).
+This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 [Rondin2010].
+The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (fig.
+12) shows a clear advantage of using FNDs in suspension for sensing.
   • ODMR: No significant shifts or changes in resonance frequency were detected, suggesting limited interaction between Gd3+ ions and the electronic spin states of the NV centers under the tested conditions (Fig.
-14.a).
+10.a).
 These findings demonstrate the potential of NV centers for detecting paramagnetic species like Gd3+.
 The sensitivity saturates at higher concentrations due to the finite interaction volume and saturation of surface sites for Gd3+ binding.
-The T1-relaxometry provides a robust method for quantifying Gd3+ in biological and environmental contexts.
-However, the effect on charge state dynamics also imply that conclusions from T1-relaxometry in otherwise uncharacterized environments should be made with caution, and at best complemented with fluorescence spectrometry.
-For completeness, we observe that by adding Gd3+ to a concentration of 0.1 M, the pH value in the solution increases slightly (by of order 1-2 pH units) which is consistent with our previous observations on the effect of pH on the fluorescence of FNDs..
-V.3 Temperature Sensing
+The fluorescence relaxometry provides a robust method for quantifying Gd3+ in biological and environmental contexts.
+However, the effect on charge state dynamics also imply that conclusions from fluorescence relaxometry in otherwise uncharacterized environments should be made with caution, and at best complemented with fluorescence spectrometry.
+For completeness, we observe that by adding Gd3+ to a concentration of 0.1 M, the pH value in the solution increases slightly (by of order 1-2 pH units) which is consistent with our previous observations on the effect of pH on the fluorescence of FNDs.
+IV.3 Temperature Sensing
 The temperature response of NV centers was evaluated under three conditions: FNDs on glass substrates with no NIR laser irradiation, trapped FNDs with no temperature increase, and trapped FNDs with temperature increases of +1 K and +2 K.
-[…]
-  • T1 Relaxation: No significant changes were detected in T1 values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig. 14.i).
-  • Optical Spectrum: No significant changes in the NV0/NV- fluorescence intensity ratio were observed with temperature, suggesting that charge state dynamics are minimally affected by thermal fluctuations (Fig.
-14.f).
+  • ODMR: A temperature-dependent shift in the resonance frequency was observed, with lower frequencies at higher temperatures (Fig.
+10.c).
+This shift is attributed to the well-known temperature dependence of the NV center’s zero-field splitting parameter $D$, which decreases at a rate of approximately $74~\text{kHz/K}$ [Acosta2010].
+  • Fluorescence relaxation: No significant changes were detected in $T_{\rm relax}$ values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig.
+10.i).
+  • Optical Spectrum: No significant changes in $\sigma_{0-}$ were observed with temperature, suggesting that charge state dynamics are minimally affected by thermal fluctuations (Fig. 10.f).
 ODMR’s temperature sensitivity highlights its utility for high-resolution nanoscale thermometry.
-The absence of significant T1 or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements
+The absence of significant $T_{\rm relax}$ or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements
+V Discussion
+The shift in the ODMR resonance frequency with increasing NIR power, as shown in Fig.
+4, Fig.
+6, and Fig.
+7a), suggests an increase in temperature within the FNDs.
+This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately 74-74 kHz/K [Acosta2010, Toyli2012].
+However, the extent of the thermal effect is significantly influenced by the surrounding medium’s properties.
+Media with higher thermal conductivity and heat capacity, such as water, dissipate heat more efficiently than air, resulting in smaller temperature increases in the FNDs.
+This is supported by our temperature simulations, which show minimal temperature rise in water compared to air during NIR illumination.
+Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts with increasing NIR power, as shown in Fig.
+6 and Fig.
+7.
+In addition to thermal effects, the NIR laser alters the charge-state dynamics of the NV centers.
+Our simulations of charge-state dynamics indicate that at low NIR power levels, there is a slight increase in the NV- population due to enhanced recombination processes facilitated by NIR photons.
+However, as the NIR power increases beyond a threshold (approximately 10 mW with our parameters), the NV- population begins to decrease while the NV0 population increases.
+This inversion is attributed to the saturation of recombination processes and increased ionization rates at higher NIR powers.
+The total fluorescence intensity decreases with increasing NIR power, as shown in Fig.
+9.
+This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers [Chapman2010].
+The increase in $\sigma_{0-}$ with higher NIR power, depicted in Fig.
+6 and Fig.
+7, further confirms the shift in charge state distribution.
+Additionally, the increased NV0 background fluorescence dilutes the ODMR contrast since NV0 centers do not exhibit spin-dependent fluorescence variations.
+This is reflected in the decreased ODMR contrast observed in our measurements (Fig.
+4).
+Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{\rm relax}$ times [Aslam2013, Goudarzi2019].
+Our results indicate that although both thermal effects and charge-state dynamics significantly impact the properties of the NV centers under NIR illumination, their relative contributions depend on the environmental conditions.
+In FNDs dried on glass substrates, the pronounced ODMR frequency shifts suggest that thermal effects dominate due to inefficient heat dissipation in air.
+In contrast, for FNDs trapped in water, the minimal temperature increase reduces thermal effects, making charge-state dynamics more prominent in influencing the NV centers’ behavior.
+The increase in the $\sigma_{0-}$ ratio and the decrease of fluorescence relaxometry times with NIR power without significant ODMR frequency shifts (Fig.
+6) indicate that charge-state transitions are the primary mechanism affecting the fluorescence and spin properties in aqueous environments.
+While thermal effects and charge-state dynamics provide a comprehensive explanation for our observations, alternative mechanisms should be considered.
+Multi-photon absorption processes could directly excite the NV centers without significant heating, yet, the relationship between ODMR frequency shift and NIR power (Fig.
+6) is contrary to this explanation.
+Moreover, the significant differences in the effects of NIR irradiation between FNDs in air and water emphasize the critical role of the surrounding medium.
+The statistical analysis using ANOVA (Table 1) confirms that the differences observed in FNDs dried on glass are statistically significant, while those in FNDs trapped in water are not.
+In our opinion, the data presented in this work establish a clear hierarchy of influences: (i) the thermal conductivity and heat capacity of the surrounding medium dictate the magnitude of the NIR-induced temperature rise and therefore the thermally induced shifts in ODMR frequency and the shortening of $T_{\rm relax}$; (ii) the NIR photon flux sets the rate of photo-ionisation and recombination, which in turn controls the NV0/NV- balance and the fluorescence contrast.
+Optical trapping itself plays a mainly mechanical role-holding a single FND at a fixed position—while potential effects on spin and charge-state dynamics depend on the wavelength of the trapping laser.
+This distinction is important for two reasons.
+First, it tells experimentalists which knob to turn: to suppress thermal artefacts one changes the medium (e.g.
+buffer composition, flow, heat sinking) rather than the trap geometry and wavelength.
+Second, it generalises our findings to intracellular sensing, where the cytoplasm’s thermal properties resemble those of water: the modest $\lvert\Delta f\rvert<2$ MHz and unchanged $T_{\rm relax}$ observed in vitro set an upper bound on in-cell NIR heating, validating the use of trapped FNDs for quantitative biosensing.
+The results presented in Section IV confirm that optical trapping of FNDs provide a valuable add-on that allows to enhance the application potential of biosensing with NV-centers in nanodiamonds.
+In particular, Fig.
+10a-c) demonstrates that the frequency shift in ODMR is a good stand-alone measure for a temperature shift also for optically trapped FNDs, while parts f) and i) illustrate that changes in the surrounding temperature within a range that is relevant for biological systems does not appear to alter the readings provided by the fluorescence intensity ratio $\sigma_{0-}$ or the relaxation time, $T_{\rm relax}$.
 VI Conclusion
-This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive understanding of the interplay between thermal effects, charge state dynamics, and their combined influence on sensing capabilities.
-By employing an integrated experimental approach combining photoluminescence spectroscopy, optically detected magnetic resonance (ODMR), and T1 relaxometry, this study elucidates how varying NIR laser powers impact the fluorescence, spin properties, and stability of NV centers in diverse environmental conditions.
-A key finding is the dual role of NIR irradiation: in dry environments, thermal effects dominate, resulting in significant shifts in ODMR frequencies and reductions in T1 relaxation times due to localized heating.
+This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive understanding of the interplay between thermal effects, charge-state dynamics, and their combined influence on sensing capabilities.
+By employing an integrated experimental approach combining photoluminescence spectroscopy, optically detected magnetic resonance (ODMR), and fluorescence relaxometry, this study elucidates how varying NIR laser powers impact the fluorescence, spin properties, and stability of NV centers in diverse environmental conditions.
+A key finding is the dual role of NIR irradiation: in dry environments, thermal effects dominate, resulting in significant shifts in ODMR frequencies and reductions in relaxation times due to localized heating.
 In contrast, for FNDs that are fully submerged in aqueous environments, the surrounding water appears to mitigate these thermal effects through efficient heat dissipation, allowing charge state dynamics to take precedence.
-The increased NV-, NV0 fluorescence intensity ratio and its implications for sensing applications underscore the critical role of the surrounding medium in tailoring NV center behavior.
+The increased $\sigma_{0-}$ ratio and its implications for sensing applications underscore the critical role of the surrounding medium in tailoring NV center behavior.
 The study also demonstrates the sensitivity of optically trapped FNDs to pH variations, temperature fluctuations, and paramagnetic species, establishing these nanodiamonds as versatile tools for nanoscale sensing.
-The results have profound implications for the development of robust NV-based sensing protocols.
 The findings indicate that optimizing NIR laser parameters, including power levels and pulsing sequences, can significantly enhance sensor performance while minimizing adverse effects such as fluorescence quenching and spin relaxation.
@@
 Theoretical modeling, while providing valuable insights into charge state dynamics and thermal effects, must be expanded to include the complexities of real-world applications.
-Future work could explore the integration of complementary sensing techniques, such as fluorescence lifetime imaging or advanced magnetic resonance methods, to provide a more holistic understanding of NV center behavior under varying conditions.
-Furthermore, efforts to miniaturize and automate the experimental setup, enabling in vivo applications, would be a critical step toward translating this research into practical diagnostic tools.
-In conclusion, this study lays a strong foundation for the continued development of NV-based sensing technologies, bridging fundamental insights with application-driven innovations.
-By addressing the remaining uncertainties and refining experimental protocols, FNDs with NV centers have the potential to become transformative tools in nanoscale sensing, advancing fields ranging from cellular biology to nanoscale physics.
-Acknowledgements.
+Future work could explore the integration of complementary sensing techniques, such as fluorescence lifetime imaging or advanced magnetic resonance methods, to provide a a real-time sensing protocol inside cells by using optically trap FNDs in the cellular medium.
+Furthermore, efforts to miniaturize and automate the experimental setup, enabling in vivo applications, would be a critical step in translating this research into practical diagnostic tools.
+acknowledgments
 This work was supported by the Independent Research Fund Denmark (grant no.
@@
 We acknowledge discussions with Maabur Sow and Fedor Jelezko, supported by ERC through HyperQ Project (SyG 856432).
-We further acknowledge insights provided through the student project work of Kristian Lambertsen and Nicolai Kongstad, including the sketch of the experimental setup, Fig.
+We further acknowledge insights provided through the student project work of Kristian Lambertsen and Nicolai Kongstad, as well as Suyash Amzare, including an original sketch of the experimental setup providing the basis for Fig.
 2.
 appendix
-As a varying range of values for the rate constants used in the numerical model exist in literature [23, 32], we provide the values used in the simulations presented here.
+We specify in Table 2 all part numbers in the optical setup, Fig. 2a).
+Table 2: Optical setup component
+Ref. Number Component Model
+1 IR laser Cobolt - Rumba 3000
+2 IR polarizer ThorLabs Multi-order polarizer
+3 IR beam dump ThorLabs BTC30
+4 IR PBS ThorLabs CCM1-PBS253/M
+5 Lens ThorLabs LA1708-B-ML
+6 Acousto-Optic Deflector AA Optoelectronics MT200 - A0.5 - 1064nm
+7 Iris Aperture ThorLabs ID12Z/M
+8 IR disk ThorLabs VRC4CPT
+9 Beam expander
+10 Right-angle prism ThorLabs DFM1/M
+11 Green laser Verdi G-Series - Coherent
+12 Optical isolator EOT 04-532-00113
+13 Green polarizer ThorLabs Multi-order polarizer
+14 Green PBS ThorLabs VA5-532/M
+15 Lens ThorLabs LA1131 - A - ML
+16 Acousto-Optic Deflector Brimrose TEM-110-25-530
+17 Iris Aperture ThorLabs SM1D12D
+18 Lens ThorLabs LA1131 - A - M
+19 Iris Aperture ThorLabs P30D
+20 Lens ThorLabs LA1509 - A
+21 Right-angle prism ThorLabs DFM1/M
+22 Lens ThorLabs LA1484 - A
+23 Filters Bightline Basic 460/60 - ThorLabs MF630-69 Emission filter
+24 Camera Teledyne Prime 95B
+25 Filter ThorLabs NE20A
+26 Filters Bightline Basic FF01-630/52 - ThorLabs FELH 0550
+27 Objective lens Olympus PLN 10X
+28 Single Photo counting module Excelitas SPCM-AQRM-13-FC
+29 Iris aperture ThorLabs SM1D12D
+30 Objective lens Nikon Plan Apo 100X
+31 Microwave Antenna Copper wire
+32 Microwave Generator Agilent N5181A
+33 Condensor Mitutuyo Plan Apo 100X
+34 Quadrant photoelectric detector ThorLabs PDQ80A
+35 White lamp
+36 Spectrometer Ocean Optics QE Pro
+In literature, a varying range of values for the rate constants used in the numerical model corresponding to Fig.
+8 exist.
+We applied the relevant values from Ref.
+[Qian2022] and combined them with parameters from [Meirzada2018] for the four ionization and recombination rates as these authors use the same laser wavelengths as us in their protocol.
+Below, in Table 3, we provide the values used in the simulations of Section III.4.
 Table 3: Parameters used in the simulation
@@
 $K_{\mathrm{NV}^{0}}^{(e)}$ 18 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ NV0 excitation rate
-$\tau_{\mathrm{NV}^{-}}$ $12\times 10^{-3}$ $\mu\mathrm{s}$ Excited-state lifetime (12 ns)
-$\tau_{\mathrm{NV}^{0}}$ $20\times 10^{-3}$ $\mu\mathrm{s}$ Excited-state lifetime (20 ns)
+$\tau_{\mathrm{NV}^{-}}$ $13\times 10^{-3}$ $\mu\mathrm{s}$ Excited-state lifetime (13 ns)
+$\tau_{\mathrm{NV}^{0}}$ $19\times 10^{-3}$ $\mu\mathrm{s}$ Excited-state lifetime (19 ns)
 $K_{\mathrm{NV}^{-}}^{(f)}$ $1/\tau_{\mathrm{NV}^{-}}$ $(\mu\mathrm{s})^{-1}$ NV- fluorescence rate
@@
 $K_{35}$ 7.9 $(\mu\mathrm{s})^{-1}$ NV- ES ms=0 $\to$ singlet
-$K_{45}$ 45 $(\mu\mathrm{s})^{-1}$ NV- ES ms=±$\pm 1$to singlet
+[…]
 $K_{56}$ 1000 $(\mu\mathrm{s})^{-1}$ Singlet ES $\to$ GS
 $K_{61}$ 6.5 $(\mu\mathrm{s})^{-1}$ Singlet GS $\to$ NV- GS ms=0
-$K_{62}$ 0.1 $(\mu\mathrm{s})^{-1}$ Singlet GS $\to$ NV- GS ms=±$\pm 1$
+[…]
 $K_{iG}$ 852 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Ionization rate (green laser)

### v3

@@
 Abstract
-Biosensing based on optically trapped fluorescent nanodiamonds is an intriguing research direction potentially allowing to resolve biochemical processes inside living cells.
+Biosensing based on optically trapped fluorescent nanodiamonds potentially allows to resolve biochemical processes inside living cells at a desired intracellular location.
 Towards this goal, we investigate near infrared (NIR) laser irradiation at 1064 nm on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers.
-The NIR wavelength is a popular choice for optical trapping in biological samples because of its low absorption in bio-samples.
-By conducting comprehensive experiments, we aim to understand how NIR exposure influences the fluorescence and sensing properties of FNDs and to determine the potential implications for the use of FNDs in various sensing applications.
-The experimental setup involved exposing FNDs to varying intensities of NIR laser light and analyzing the resultant changes in their optical and physical properties.
+The 1064 nm NIR wavelength is a popular choice for optical trapping because of its low absorption in bio-samples.
+By conducting comprehensive experiments, we aim to understand if and how NIR exposure influences the fluorescence and sensing capabilities of FNDs and to determine the potential implications for the use of FNDs in various sensing applications.
+Our experiments exposed FNDs to varying intensities of NIR laser light while carefully monitoring their optical and magnetic properties.
 Key measurements included all-optical fluorescence relaxation, optical spectroscopy, and optically detected magnetic resonance (ODMR) spectra.
-The findings reveal how increased NIR laser power correlates with alterations in ODMR central frequency but also that charge state dynamics under NIR irradiation of NV centers plays a role.
-We suggest protocols with NIR and green light that mitigate the effect of NIR, and demonstrate that FND biosensing works well with such a protocol.
+The findings reveal how increased NIR laser power correlates with alterations in ODMR central frequency but also that charge state dynamics under NIR irradiation of NV centers play a role.
+We demonstrate that FND biosensing works well with a protocol involving both NIR and green light, while mitigating the effect of NIR.
+keywords:
 Nanodiamond , Optical trapping , Temperature sensing , NV-centers
-preprint: APS/123-QED
-I Introduction
-Figure 1: Graphical abstract illustrating the sensing modalities explored to investigate effects of NIR laser on NV center for sensing.
-Temperature, abundance of paramagnetic species, and pH, are altered while optical spectroscopy, ODMR spectra, and fluorescence polarization relaxometry is measured from optically trapped nanodiamonds.
-Fluorescent nanodiamonds (FNDs) hosting negatively-charged nitrogen-vacancy (NV-) centers combine bright, photostable fluorescence with quantum-coherent electron-spin states, making them uniquely suited for bio-sensing and bio-imaging at the nanoscale [Schirhagl2014, Liu2021].
-Because NV spins can be optically polarised and read out at room temperature, changes in their spin-state populations translate directly into optical signals, enabling quantitative sensing of magnetic fields, temperature and local electro-chemical potentials [Doherty2013, Rondin2014].
+journal: Diamond and related materials
+[…]
+[…]
+{graphicalabstract}[Uncaptioned image]
+{highlights}
+Demonstration of multi-modal NV-center based sensing with optically trapped nanodiamonds
+Exploring the effect of a 1064 nm NIR trapping laser on the NV-center readouts
+1 Introduction
+Fluorescent nanodiamonds (FNDs) hosting negatively-charged nitrogen-vacancy (NV-) centers combine bright, photostable fluorescence with quantum-coherent electron-spin states, making them uniquely suited for bio-sensing and bio-imaging at the nanoscale Schirhagl et al.
+(2014); Hsiao et al.
+(2016).
+Because NV spins can be optically polarised and read out at room temperature, changes in their spin-state populations translate directly into optical signals, enabling quantitative sensing of magnetic fields, temperature and local electro-chemical potentials Doherty et al.
+(2013); Rondin et al.
+(2014).
 The ultimate performance of those sensors, however, is governed by the spin- and charge-state photodynamics of the NV centers themselves; understanding how an experimental control parameter perturbs those dynamics is therefore a prerequisite for accurate, high-sensitivity measurements.
-The NV center is an electronic spin-1 system with a ground-state zero-field splitting (ZFS) $D\approx 2.87~\mathrm{GHz}$ that separates the $m_{s}=0$ and $m_{s}=\pm 1$ sub-levels.
-External magnetic fields lift the degeneracy of the $m_{s}=\pm 1$ levels through the Zeeman effect, while lattice strain, temperature and electric fields shift the ZFS and the optical transition energies via the Stark effect [Maze2008, Dolde2011].
+The negatively charged NV center is an electronic spin-1 system with a ground-state zero-field splitting (ZFS) $D\approx 2.87~\mathrm{GHz}$ that separates the $m_{s}=0$ and $m_{s}=\pm 1$ sub-levels.
+External magnetic fields lift the degeneracy of the $m_{s}=\pm 1$ levels through the Zeeman effect, while lattice strain, temperature and electric fields shift the ZFS and the optical transition energies via the Stark effect Maze et al.
+(2008); Dolde et al.
+(2011).
 In addition, the relative abundance of the neutral (NV0) and negative (NV-) charge states—and thus the fluorescence spectrum and intensity—depends sensitively on the local Fermi level and on photo-ionisation/recombination pathways.
-These intertwined dependencies let NV centers act as multi-modal nano-probes, provided that the separate interaction between all external controls (laser beams, microwave fields, etc.) and the FNDs are themselves well understood and do not compromise sensor performance.
+These intertwined dependencies let NV centers act as multi-modal nano-probes of their local environment, provided that the separate interaction between all external controls (laser beams, microwave fields, etc.) and the FNDs are themselves well understood and do not compromise sensor performance.
 A particularly powerful external control is optical trapping.
-Near-infra-red (NIR) optical tweezers can confine—and manoeuvre—sub-100  nm diamonds inside living cells or complex micro-fluidic environments, offering spatially resolved sensing with a resolution set by the nanoparticle size rather than by the optical diffraction limit [Neukirch2013].
-Operating the trap at 1064  nm keeps tissue absorption low and mitigates photo-damage [Svoboda1994], yet the same intense NIR field may also heat the nanodiamond, alter NV charge-state kinetics or modify spin-lattice relaxation.
-Earlier work on bulk diamond or levitated nanodiamonds indeed reported substantial NIR-induced changes in fluorescence and $T_{\rm relax}$[Aslam2013, Ji2016], but a clear picture for FNDs dispersed in aqueous media—arguably the configuration of greatest practical relevance-is missing.
-Here, as illustrated by Fig.1, we address that gap.
-Using freely diffusing and optically trapped FNDs in water and in biologically relevant buffer solutions, we quantify how a 1064  nm focused trapping beam (30–60  mW average power) affects NV optical and spin properties.
-We employ tree complementary read-outs:
-  1. 1.
-    photoluminescence (PL) spectroscopy to monitor charge-state redistribution [Geiselmann2013];
-  2. 2.
-    optically detected magnetic resonance (ODMR) to track shifts in the resonance frequencies with kHz precision [Gruber1997, Sow2020]; and
-  3. 3.
-    fluorescence relaxometry to probe magnetic-noise spectra—e.g. from free radicals or $\mathrm{Gd}^{3+}$ ions—at the NV transition frequency [Hall2016].
-Together, these observables allow us to disentangle three mechanisms that NIR light can induce:
-  1. 1.
-    Photothermal heating of the FND, shifting the ZFS via the temperature coefficient $dD/dT\approx-74~\mathrm{kHz\,K^{-1}}$[Acosta2010];
-  2. 2.
-    Power-dependent ionisation/recombination that re-weights NV- and NV0, thereby changing both ODMR spectral properties and PL brightness [Aslam2013, Ji2016];
-  3. 3.
-    Modified surface sensitivity, whereby the NIR field—through either heating or altered charge state—amplifies or suppresses the NV response to environmental parameters such as pH or paramagnetic species concentration [Fu2010].
-We first benchmark these effects in air-dried FNDs, then in trapped FNDs immersed in solutions of controlled pH and temperature, and finally in the presence of a paramagnetic relaxant ($\mathrm{Gd}^{3+}$).
-By comparing experiments with and without the NIR trap, we establish quantitative limits on NIR-induced artefacts and delineate operating regimes where accurate, spatially resolved biosensing is possible.
-II Methods
-II.1 Sample Preparation
-Fluorescent FNDs with a nominal diameter of 120 nm (Sigma-Aldrich, 798088) were used for most of the experiments reported in this article if not stated otherwise.
-They possess a significant number of nitrogen-vacancy (NV) centers due to a high concentration of 3ppm.
-To prepare the FND suspension, 1 µL of the commercial stock solution of FNDs was diluted in 300-700 µL of deionized water depending on the desired concentration and before sonicating for 10 minutes.
-Next, the FND suspension was deposited on a petri dish (Mattek, part no P35G-1.5-10-C.), thoroughly cleaned to remove any contaminants that could interfere with the experiments and plasma ionised in order to make the surface hydrophilic, ensuring an even distribution of FNDs on the surface.
+Near-infra-red (NIR) optical tweezers can confine and manoeuvre sub-100  nm diamonds inside living cells or complex micro-fluidic environments, offering spatially resolved sensing with a resolution set by the nanoparticle size rather than by the optical diffraction limit Horowitz et al.
+(2012).
+Operating the trap at 1064  nm keeps tissue absorption low and mitigates photo-damage Svoboda and Block (1994), yet the same intense NIR field may also heat the nanodiamond, alter NV charge-state kinetics or modify spin-lattice relaxation Lai et al.
+(2013); Geiselmann et al.
+(2013); Ji and Dutt (2016).
+Earlier work on bulk diamond or levitated nanodiamonds indeed reported substantial NIR-induced changes in fluorescence and the spin relaxation $T_{\rm relax}$Neukirch et al.
+(2013); Ji and Dutt (2016), but a clear picture for FNDs dispersed in aqueous media—arguably the configuration of greatest practical relevance-is missing.
+Here, as illustrated by the graphical abstract, we systematically address that gap.
+Using freely diffusing and optically trapped FNDs in water and in biologically relevant buffer solutions, we quantify how a 1064 nm focused trapping beam with an average power between 30–60 mW focused to a spot with an Airy diameter less than 450 nm affects NV optical and spin properties.
+We employ three complementary read-outs, each indicative of a particular NV property: i) photoluminescence (PL) spectroscopy to monitor charge-state redistribution between NV0 and NV-Aslam et al.
+(2013); Geiselmann et al.
+(2013); Sow et al.
+(2020); ii) optically detected magnetic resonance (ODMR) at zero bias field to track shifts in the resonance frequencies with kHz precision which we discuss as a measure of temperature change  Gruber et al.
+(1997); Acosta et al.
+(2010); Doherty et al.
+(2014); Hall et al.
+(2016); and iii) fluorescence relaxometry to probe magnetic-noise spectra—e.g.
+from free radicals or $\mathrm{Gd}^{3+}$ ions—at the NV transition frequency Tetienne et al.
+(2013); Mzyk et al.
+(2022).
+Together, these observables allow us to address and potentially disentangle three mechanisms that NIR light has been suggested to induce, namely a) Photothermal heating of the FND, shifting the ZFS via the temperature coefficient $dD/dT\approx-74~\mathrm{kHz\,K^{-1}}$Acosta et al.
+(2010); Lai et al.
+(2013); b) Power-dependent ionisation/recombination that re-weights NV- and NV0, thereby changing both ODMR spectral properties and PL brightness Aslam et al.
+(2013); Ji and Dutt (2016); and c) Modified surface sensitivity, whereby the NIR field—through either heating or altered charge state—amplifies or suppresses the NV response to environmental parameters such as pH or paramagnetic species concentration Fu et al.
+(2010).
+Initially, we compare these effects in air-dried FNDs, then in trapped FNDs immersed in an aqueous solution.
+Then, to address potential impact of NIR on sensing capabilities of the FNDs, we measure on trapped FNDs in solutions of controlled pH and temperature, and finally in the presence of a paramagnetic relaxant ($\mathrm{Gd}^{3+}$).
+Thereby, we establish quantitative limits on NIR-induced artefacts and delineate operating regimes where accurate, spatially resolved biosensing is possible.
+2 Methods
+2.1 Sample Preparation
+Fluorescent FNDs with a nominal diameter of 120 nm (Sigma-Aldrich, 798088) and an NV concentration of 3ppm were used for most of the experiments reported in this article if not stated otherwise.
+To prepare the FND suspension, 1 µL of the commercial stock solution was diluted in 300-700 µL of deionized water depending on the desired concentration and before sonicating for 10 minutes.
+Next, the petri dish (Mattek, part no P35G-1.5-10-C) was prepared before deposition of FNDs.
+For this, the petri dish was thoroughly cleaned and plasma ionised in order to make the surface hydrophilic, ensuring an even distribution of FNDs on the surface.
 Plasma treatment was applied for 30 seconds in a simple custom-device with a closed chamber in PMMA and under ambient conditions.
 Plasma was generated with a BD-20AV high-frequency generator (230 V, ETP, IL, USA) with an electrode positioned close to the petri dish surface.
-The plasma treatment modifies the surface energy of the petri dish by breaking chemical bonds on the surface and introducing polar functional groups, thereby inducing the hydrophilic state [Al-Azzam2022].
-Approximately 3 µL of the FND suspension was then carefully deposited onto the clean petri dish using a micropipette.
-The suspension was gently spread across the surface before allowing it to dry.
-The drying process took place at room temperature to minimize any potential alterations to the FND properties potentially caused by excessive heating or cooling.
-After the solution had completely evaporated, the sample was ready for the optical experiments described in the following.
+Approximately 3 µL of the FND suspension was then carefully deposited onto the clean petri dish using a micropipette and the sample was left to dry at room temperature.
 The preparation of FNDs for experiments in the liquid environment also starts with the dilution steps and cleaning of the petri dish, but to maintain the FNDs in solution, the petri dish was not plasma treated and the solution was not left for evaporation.
-FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments [Niora2024].
-The FNDs (Sigma-Aldrich 1003626896) were prepared following the same protocol.
+FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments Niora et al.
+(2024); Mzyk et al.
+(2025, 2026).
+The 70 nm FNDs (Sigma-Aldrich 1003626896) were prepared following the same protocol.
 A suspension of FNDs was created by diluting 1 µL of the commercial stock solution in 1000 µL of deionized water, followed by sonication for 10 minutes to ensure a homogeneous dispersion.
 Approximately 50 µL of the suspension was deposited in a sterile Petri dish that was freshly unwrapped.
-II.2 Optical Setup
+2.2 Optical Setup
 The dried FND sample was placed inside a mini-incubator (pecon XS 2000) in an optical setup based on a customized Thorlabs OTKB/M modular optical tweezer system.
-Customization included a different objective and long-distance condenser lens.
-Vendors and other details are listed in Table 2 in the Appendix.
-The setup is constructed around an inverted microscope configuration, equipped with a high-resolution camera that capture wide-field bright-field images of the FND, and two lasers for the simultaneous illumination and manipulation of the FNDs.
-A green laser with a wavelength of 532 nm was utilized for the excitation of the NV centers, enabling the initialization and readout of their spin states.
-The second laser is operating in the near-infrared (NIR) at a wavelength of 1064 nm and used for FND trapping.
-This wavelength falls within the biological transparency window, which minimizes absorption and reduces the risk of significant heating effects in biological samples.
-The two lasers were aligned and their beams combined using a dichroic mirror, ensuring that they shared the same optical path before being focused onto the sample through the microscope objective, as detailed in Fig.
-2a).
+Customization included a different trapping laser, and different objective and long-distance condenser lens.
+Vendors and other details are provided in the Appendix, Fig.
+9 and Table 2.
+The setup is constructed around an inverted microscope configuration, equipped with white light illumination and a high-resolution camera that capture wide-field bright-field images of the sample, and two lasers for the simultaneous illumination and manipulation of the FNDs.
+A 532 nm laser was utilized for the excitation of the NV centers, while a 1064 nm near-infrared (NIR) laser was used for FND trapping.
+The two lasers were aligned and their beams combined using a dichroic mirror and collectively focused onto the sample through the microscope objective, as detailed in Fig.
+1, i.e., a single objective setup.
 By trapping an FND and then finetuning the position of the green laser to maximize the signal on the photon counter, we ensured the same focal point for both the green and NIR lasers.
-Additionally, acousto-optic deflectors were used to precisely control the timing and intensity of the laser beams by using Qudi software, enabling fine-tuned manipulation of the FNDs.
+Additionally, acousto-optic deflectors, controlled via a Qudi Binder et al.
+(2017) interface, were used to precisely control the timing and intensity of the laser beams.
 As illustrated in Fig.
-2a), several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection.
-A motorized flip-mirror ensures the ability to rapidly shift between camera detection and either spectrometer or photon count detection.
-The fluorescence spectrum of the FNDs is recorded and analyzed by a fiber coupled spectrometer while alternatively photon counts are measured with a fiber coupled avalanche photo detector.
-In addition, to measure the CW ODMR spectrum of FNDs, a signal generator connected to an amplifier is connected to a single-turn copper loop antenna (5 mm) placed approximately 1 mm above the sample.
-When the flip-mirror is positioned towards the camera, the latter captures high-resolution images of the FNDs illuminated during the experiments in real time.
-The combination of different detection methods thus made it possible to continuously monitor the position of the FNDs, determine whether they were isolated or in aggregates, and analyse their photoluminescence properties throughout the experimental process.
-Figure 2: a) Illustration of the optical setup with the most important lasers, optical lenses, filters and equipment.
-Details of the components are specified in Table 2 in the appendix.
-b) Laser sequence for fluorescence relaxometry measurement.
+1, several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection.
+A motorized flip-mirror allows to rapidly shift between camera detection and either spectrometer or photon count detection, for recording of FND fluorescence.
+In addition, to measure the CW ODMR spectrum of FNDs, a signal generator together with an amplifier are connected to a single-turn copper loop antenna (5 mm) placed approximately 1 mm above the sample.
+Figure 1: Essential elements of the optical setup.
+NV centers are excited with 532 nm green light, and nanodiamonds are optically trapped with the 1064 nm NIR light.
+Detection modalities include bright-field, wide-field imaging on a camera as well as photon counting and fluorescence spectroscopy.
+Zoom on the sample shows a) the situation with NIR and green light focused on an FND on glass, in dry conditions.
+Similarly b) shows the same but with added water while c) illustrates the FND being optically trapped in aqueous medium.
+d) Laser sequence for fluorescence relaxometry measurement.
 In green the 532 nm laser and in red the 1064 nm laser.
 Pulse-widths are $t_{\rm trap}=20\mu s$, $t_{ini}=35\mu s$, $t_{RO}=15\mu s$ while the variable waiting time $\tau$ allows to record the relaxometry at times spanning from $7.2\mu s$ to $720\mu s$.
-II.3 FND Trapping
+2.3 FND Trapping
 All trapping experiments were performed in a fluid chamber containing 50 µL of the FND aqueous suspension unless stated otherwise.
-The NIR laser was utilized for trapping the FNDs in the focused beam [Neuman2004, Gennerich2017, Jones2015].
+The NIR laser was utilized for trapping the FNDs in the focused beam Neuman and Block (2004); Jones et al.
+(2015); Gennerich (2017).
+FND trapping was confirmed through standard analysis of histograms and power spectral density of the positions of the FND Berg-Sørensen and Flyvbjerg (2004), with positions recorded by back focal plane interferometry on a quadrant photodiode Allersma et al.
+(1998).
+Examples are shown in the Supplementary information.
 For relaxometry, the temporal sequence of the green and trapping lasers, as illustrated in Fig.
-2b), avoids temporal overlap between the two lasers while maintaining the FND stably trapped [Russell2018a].
+1d), avoids temporal overlap between the two lasers while maintaining the FND in the trap Russell et al.
+(2018).
 Strictly speaking, this protocol does not constitute a conventional spin-relaxometry measurement.
-With an all-optical protocol, the method inherently involves potential inaccuracies due to photoionization induced by the green laser itself, thus affecting the NV- charge states.
-To address this limitation, measurement protocols incorporating microwave pulses explicitly designed to disentangle spin dynamics from photoionization effects have been developed.
-Examples include the microwave $\pi$-pulse method introduced by Ref.
-[Jarmola2023], as well as alternative sequences described in Ref.
-[Cardoso2023].
 We hence will refer to the measured decay time as $T_{\rm relax}$ throughout the remainder of this paper.
-The green laser was pulsed with initialization time $t_{ini}$ of 35 µs, and read-out time $t_{RO}$ of 15 µs, for the readout of the NV center fluorescence, while the NIR laser was on to trap the FNDs in sequences with duration $t_{\rm trap}$ of 20 µs.
-II.4 Sensing Experiments
+The green laser was pulsed with NV initialization time $t_{ini}$ of 35 µs and read-out time $t_{RO}$ of 15 µs, while the NIR laser was turned on with duration $t_{\rm trap}$ of 20 µs to trap the FNDs.
+2.4 Sensing Experiments
 For the experiments detecting paramagnetic species, a gadolinium solution (0.45 M in acetate) was gradually added to the FND’s environment.
-The concentrations were incrementally adjusted, and measurements were performed after each addition to evaluate the changes in the NV center properties.
+The concentrations were incrementally adjusted, and measurements on FNDs were performed after each addition to evaluate the changes in the NV center properties.
 For the pH sensing experiments, an initial 0.1 M NaOH(aq) solution at pH 13 was prepared.
 The pH was then progressively decreased by adding controlled volumes of a 1 M HCl solution.
-After each pH adjustment, the properties of the NV centers were measured while the FNDs were maintained in the optical trap at 60 mW.
+After each pH adjustment, the properties of the NV centers were measured while the FNDs were maintained in the optical trap with a power of 60 mW, corresponding to an intensity of order $10^{7}$ W/cm2.
 For the temperature sensing experiments, the medium containing the FNDs was initially maintained at 37 C using an incubator.
@@
 For the sensing of paramagnetic species and pH, changes to the surrounding medium was involved.
-To ensure that the FND remained trapped despite disturbances in the medium, the CW NIR laser power was temporarily increased to 100 mW while liquid was added.
+To ensure that the FND remained trapped despite disturbances in the medium, the CW NIR laser power was temporarily increased to around 100 mW while liquid was added.
 Subsequently, the laser power was reduced back to 60 mW for measurements with the new conditions, and the same FND.
-III Results
-We begin by presenting the main results of our characterisation experiments, starting with the examination of the general fluorescence properties of NV-centers within the FNDs through photoluminescence (PL) spectroscopy.
-This foundational assessment provides an overview of the baseline optical characteristics and helps identify any immediate changes induced by the NIR laser irradiation.
-Following the PL spectrum analysis, we present the variations in fluorescence intensity as a function of optically detected magnetic resonance (ODMR) conditions.
-This step is crucial for understanding the impact of NIR irradiation on the magnetic resonance properties of the NV centers.
-Finally, we conduct fluorescence relaxometry to investigate the spin relaxation dynamics under different NIR laser parameters.
-PL spectrum measurement and ODMR measurement were conducted with CW NIR laser irradiation.
-The experiments conducted under dry conditions primarily serve as a thermal control benchmark.
-This dry setup facilitates the understanding of thermal phenomena, such as localized heating effects, by providing a clear reference to compare with conditions involving FNDs suspended in aqueous environments.
-We acknowledge that the dry condition itself lacks direct practical application in sensing contexts.
-For each set of experiments described in the subsections below, we exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW.
-III.1 Photoluminescence Spectrum Analysis
-As the NV neutral and negative charge states have distinct optical emission spectra, we conducted PL spectroscopy of FNDs at varying NIR power to investigate its impact.
-The FNDs were either dried on glass substrates or trapped in water to examine the influence of the surrounding medium.
-We record the PL spectrum in the range between 550 nm to 745 nm with the spectrometer for each of the different values of the CW NIR laser power.
-Figure 3: Effect of CW NIR laser on dry FND emission spectrum.
-Optical emission spectrum of one FND, dry on glass, under CW NIR laser exposure at 0 mW (green), 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
-Upon CW NIR laser light exposure, our PL measurements revealed a significant drop in the total PL intensity.
-Notably, as a function of NIR laser power there was a slightly more pronounced reduction in the spectral intensity of NV- compared to NV0.
-For the fluorescence spectrum (like Fig.
-3), the curves remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30mW to 60mW (data not shown) and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water.
-To quantify the changes in relative charge state emission intensities under CW NIR laser irradiation, we determine the ratio $\sigma_{0-}$ between NV0 and NV- emission intensities by integrating over the spectral ranges for NV0 (550–620 nm) and NV- (655–745 nm), respectively.
-Compared to the analysis used in Ref.
-[Sow2020], we applied a normalization factor 19/14 to account for the difference in the relative spectral ranges:
-$\sigma_{0-}\equiv\left(\frac{\sum_{\lambda=550}^{620}I_{\lambda}}{\sum_{\lambda=655}^{745}I_{\lambda}}\right)\times\frac{19}{14}$ (1)
-For FNDs trapped with NIR powers of 30 mW and 60 mW, we observe an averaged similar emission ratios $\sigma_{0-}$ of $0.48\pm 0.09$.
-In contrast, for FNDs dried on glass, we measure an emission ratio of $0.43\pm 0.11$ without NIR irradiation that increases to $0.57\pm 0.15$ for a NIR laser power of 60 mW NIR.
-All results are displayed in Fig.
-6 parts a) and c).
-III.2 Optically Detected Magnetic Resonance Measurements
-In the next step, we elaborate on the influence of CW NIR laser light on the ODMR resonance frequency and contrast, revealing the impact of the NIR trapping laser on the NV electron spin properties.
-The ODMR spectra were acquired in continuous-wave (CW) mode with constant intensities of both the 532 nm pump and 1064 nm trapping lasers and the microwave source, while sweeping the frequency of the microwave in steps of 0.3 MHz.
-Generally and for all FNDs, we observe a redshift in the resonance frequency with increasing CW NIR laser power, as illustrated in Fig.
-4 by the spectra from one FND with 0 mW and 60 mW NIR power.
-Figure 4: Effect of CW NIR laser power on ODMR spectrum of dry FND.
-ODMR spectra of one FND under CW NIR laser exposure at 0 mW (green) and 60 mW (red) of 1064 nm laser power.
-The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed.
-The power of the 532 nm laser was 0.7 mW, measured before the objective lens, and the microwave amplifier was set to -2dB.
-For FNDs dried on glass substrates, the resonance frequency shifted significantly to the lower frequencies.
-Taking 2.87 GHz as reference value for the zero field splitting, the resonance frequency shift increased in absolute value from $0.31\pm 0.54$ MHz at 0 mW to $5.83\pm 046$ MHz at 60 mW of CW NIR power.
-In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $0.84\pm 0.21$ MHz at 30 mW and $1.58\pm 0.33$ MHz at 60 mW.
-Examples are shown in Fig.
-4 and combined results displayed in Fig.
-7.
-As before, for the similar experiment when water was added to the same FND that remained stuck on the glass slide, the ODMR spectra show a trend very similar to Fig.
-4 although the difference between 30 mW and 60 mW of NIR power is less pronounced (data not shown).
-[…]
-III.3 Relaxometry Measurements
-Last, we conducted relaxometry experiments under different NIR laser power, to extract values for the corresponding $T_{\rm relax}$.
-As above, FNDs were either dried on glass substrates or suspended in water to assess the influence of the surrounding medium on the relaxometry measurements with NIR irradiation.
-For these experiments, the all-optical pulsing sequence in Fig.
-2b) was applied.
-[…]
-Figure 5: Effect of NIR laser power on $T_{\rm relax}$ relaxation times of dry FNDs.
-$T_{\rm relax}$ relaxometry measurements for one FND dried on glass under NIR laser powers of 0 mW (green) and 60 mW (red) at 1064 nm wavelength.
-Pulsing as in Fig 2b).
-[…]
-Overall, as showed in Fig.
-6 and Fig.
-7, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts.
-Further, our results confirm that with specific NIR irradiation sequences and optimizing laser parameters, cf.
-Fig.
-2b) and Ref.
-[Russell2018a], potential detrimental effects of NIR on $T_{\rm relax}$ relaxation can be significantly mitigated.
-Figure 6: a) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs trapped inside water.
-b) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs trapped inside water.
-c) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs dry on glass.
-d) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs dry on glass.
-Data for 30 FNDs exposed to 0 mW (green), 30 mW (orange), and 60 mW (red) of 1064 nm NIR laser irradiation.
-CW NIR for ODMR and spectra, pulsed NIR as in Fig 2b) for fluorescence relaxometry.
-In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and and no or much less pronounced trend for FNDs in the optical trap in water.
-Figure 7: Summary of experimental results, showing variation for individual nanodiamonds.
-(a) ODMR resonance frequency shift, relative to the expected resonance at 2.87 GHz, (b) fluorescence intensity ratio $\sigma_{0-}$, and (c) $T_{\rm relax}$ relaxometry times as functions of NIR laser power.
-Data for 5 random FNDs dry on glass (black stars) and trapped in water (gray circles) with 1064 nm NIR laser irradiation.
-To study the significance of the observed effects, we performed an Analysis of Variance (ANOVA) test on the $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
-The ANOVA results are summarized in Table 1 (one-way Kruskal–Wallis test, $\alpha=0.05$).
-Table 1: ANOVA test results for the effect of NIR laser power on FNDs dried on glass substrates, on glass after adding water and trapped in water.
-The p-values indicate the significance of differences observed across NIR power levels for $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
-p-value
-FNDs dried on glass
-$T_{\rm relax}$ Relaxation Time <0.0001<0.0001
-ODMR Frequency Shift <0.0001<0.0001
-$\sigma_{0-}$ Ratio 0.0024
-FNDs on glass with water
-$T_{\rm relax}$ Relaxation Time $0.072$
-ODMR Frequency Shift $0.008$
-$\sigma_{0-}$ Ratio 0.201
-FNDs Trapped in Water
-$T_{\rm relax}$ Relaxation Time 0.221
-ODMR Frequency Shift 0.073
-$\sigma_{0-}$ Ratio 0.371
-For FNDs dried on glass substrates, the very low p-values (p <0.01<0.01) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant.
-This indicates that NIR irradiation has a considerable impact on the NV centers’ properties in air.
-For FNDs on glass substrates after adding water, the mean p-values ($0.01<p<0.2$) for all three parameters confirm that the differences observed with varying NIR laser powers are not always statistically significant.
-In fact, the difference between no NIR laser and NIR laser is significant but the differences between 30 mW and 60 mW power of NIR laser are not significant.
-In contrast, for FNDs trapped in water, the p-values are high (p >0.07>0.07), suggesting that the differences observed across different NIR power levels are not statistically significant.
-This reinforces the conclusion that the aqueous environment mitigates the effects of NIR irradiation on the NV centers.
-III.4 Model of Charge State Dynamics
-To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model [Qian2022], supplemented with insights from [Meirzada2018].
+2.5 Simulations - Model of Charge State Dynamics
+To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model Qian et al.
+(2022), supplemented with insights from Ref.
+Meirzada et al.
+(2018).
 This model accounts for the various optical transitions, inter-system crossings, ionization, and recombination processes that occur when NV centers are simultaneously exposed to green and NIR laser excitation.
-The NV center can exist in two charge states: the negatively charged NV- and the neutrally charged NV0.
-Each charge state has its own ground and excited electronic states, as well as inter-system crossing pathways to metastable quartet states.
-The model considers 8 states as illustrated in the schematic energy level diagram in Fig.
-8.
-The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state [Qian2022].
-The rates of change were adapted to our experimental situation with just two lasers, the green 532 nm and the NIR 1064 nm laser, following [Meirzada2018].
-Values used in this work for the rates are provided in the appendix, Table 3.
-The model is based on the following assumptions:
-  • Steady-State Approximation: The populations reach a steady state within the simulation time, allowing us to analyze the equilibrium behavior.
-  • Neglect of Higher-Order Processes: Multi-photon processes and higher excited states are neglected due to their low probabilities under the experimental conditions.
-  • Constant Rate Coefficients: The rate coefficients are assumed to be constant and independent of the local temperature and environmental factors.
-  • No External Fields: Effects of external magnetic or electric fields on the transition rates are not included.
-Figure 8: Energy level diagram of the NV center, showing the ground states (P1, P2 and P7), excited states (P3, P4 and P8), quartet states (P5 and P6), and the transitions between them.
+The model considers eight states as illustrated in the schematic energy level diagram in Fig.
+2.
+These states include the negatively charged NV- and the neutrally charged NV0, with each their ground and excited electronic states, as well as inter-system crossing pathways to metastable quartet states.
+The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state $i=1,\ldots 8$ Qian et al.
+(2022).
+The rates of change were adapted to our experimental situation with just two lasers, the green 532 nm and the NIR 1064 nm laser, following Ref.
+Meirzada et al.
+(2018).
+Values used in this work for the rates are provided in the appendix, Table 3, and rate-equations are explicitly written in the Supplementary information.
+Simulations were carried out using the solve_ivp function from the sciPy library in python.
+The model assumes that populations reach a steady state within the simulation time, and that transitions happen with constant rate coefficients, independent of local temperature and other environmental factors including external magnetic or electric fields.
+Further, the model neglects multi-photon processes as well as higher excited states as the probabilities for these are negligible under the experimental conditions.
+Figure 2: Energy level diagram of the NV center, showing the ground states $^{3}A_{2},m_{s}=0$ (denoted as 1), $^{3}A_{2},m_{s}=\pm 1$ (2) and $^{2}E$ (7), excited states $^{3}E,m_{s}=0$ (3), $^{3}E,m_{s}=\pm 1$ (4) and $^{2}A$ (8), singlet shelving states $^{1}E$ (5) and $^{1}A_{1}$ (6), and the transitions between them.
 Solid arrows represent radiative transitions, dotted arrows represent non-radiative transitions, and dashed arrows represent ionization and recombination processes.
-The simulation results reveal that at low CW NIR laser powers, the population of NV- centers increases slightly, while the NV0 population decreases.
-This is attributed to enhanced recombination processes facilitated by the NIR photons, which convert NV0 centers back to NV-.
-Specifically, up to an CW NIR power of approximately 10 mW, the NV- population increases by about 8% compared to the case without CW NIR illumination.
-However, as the CW NIR power increases beyond this threshold, the trend reverses.
-The NV- population begins to decrease, and the NV0 population increases.
-At a CW NIR power of 60 mW, the NV- population decreases by approximately 10% from its maximum value, while the NV0 population increases correspondingly.
-This inversion is due to the saturation of recombination processes and the dominance of ionization processes at higher CW NIR powers, which ionize NV- centers back to NV0.
-The total fluorescence intensity from the NV centers decreases with increasing CW NIR power (Fig.
-9).
-This decrease is primarily due to the reduced population of NV- centers and the lower fluorescence quantum yield of NV0 centers.
-The fluorescence from NV- centers dominates the emission under low CW NIR power, but as the CW NIR power increases, the contribution from NV0 centers becomes more significant and we observe a saturation in the fluorescence.
-The $\sigma_{0-}$ ratio increases with CW NIR power.
-At low CW NIR powers, this ratio is low due to the predominance of NV- fluorescence.
-As the CW NIR power increases, the ratio rises slowly, indicating a relative increase in NV0 emission.
-This change in the fluorescence ratio affects the overall spectral characteristics and can influence the detection sensitivity in applications relying on NV- fluorescence.
-The trends observed in the simulation align with the experimental results.
-The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity (Fig.
-9) and $\sigma_{0-}$ ratio (Fig.
-6a,c)) observed experimentally.
-Figure 9: Total fluorescence intensity from NV centers as a function of NIR laser power.
-(Black: Theory, Red: Experimental observations with errorbars, obtained as a statistical average over results from 10 different FNDs with $\sigma_{0-}$ within the range of $\sigma_{0-}=0.42\pm 0.04$.)
-III.5 Temperature Simulation
-To model the temperature increase induced by the NIR laser on FNDs, we simulated the heat transfer.
+2.6 Simulation - Thermal response
+To model a potential temperature increase induced by the NIR laser on FNDs, we simulated the heat transfer.
 The simulation accounts for the absorption of laser energy by the FND and the subsequent heat dissipation into the surrounding medium (water or air).
-To quantify the temperature rise produced by the 1064 nm laser on a spherical fluorescent nanodiamond (radius $R=35\;\text{nm}$), we solve Fourier’s heat‑conduction equation with an internal absorption source inside the particle and pure conduction in the surrounding infinite medium (water or air):
+To quantify the temperature rise produced by the 1064 nm laser on a spherical fluorescent nanodiamond of radius $R$), we solve Fourier’s heat‑conduction equation with an internal absorption source inside the particle and pure conduction in the surrounding infinite medium (water or air):
 $\rho_{d}c_{d}\frac{\partial T_{d}}{\partial t}$ $=\frac{1}{r^{2}}\frac{\partial}{\partial r}\!\left(k_{d}r^{2}\frac{\partial T_{d}}{\partial r}\right)+q_{\mathrm{abs}}(r),$ $0\leq r<R,$
@@
 where $P_{\mathrm{abs}}=\sigma_{\mathrm{abs}}I_{0}$ is the absorbed optical power for a beam of intensity $I_{0}$ and absorption cross‑section $\sigma_{\mathrm{abs}}$.
-The simulation is based on the following assumptions:
-  • Uniform Temperature Distribution: The FND is assumed to have a uniform temperature at any given time due to its small size.
-  • Negligible Radiation Losses: Heat loss due to thermal radiation is neglected because it is insignificant compared to conduction at the temperatures considered.
-  • Constant Thermal Properties: The thermal properties of the FND and the medium are assumed constant over the temperature range studied.
-  • Steady-State Heat Transfer Coefficient: The convective heat transfer coefficient is considered constant and based on the thermal conductivity of the medium.
-  • No Phase Change: The medium remains in the same phase (liquid for water, gas for air) throughout the simulation.
+Because of the small size of the FNDs, the simulation assumes a uniform temperature throughout the FND at any given time, it assumes insignificant heat loss due to thermal radiation, and constant thermal properties of the FND and the medium, including a constant convective heat transfer coefficient based on the thermal conductivity of the medium.
+Further, the model assumes that the medium remains in the same phase (liquid) throughout the simulation.
+3 Results
+Our initial characterisation experiments encompass ODMR and photoluminescence (PL) spectroscopy, and fluorescence relaxometry, with and without NIR irradiation.
+PL spectrum measurement and ODMR measurement were conducted with CW NIR laser irradiation while relaxometry experiments applied the pulse protocol in Fig.
+1d).
+Three conditions were investigated, either with FNDs on glass under dry conditions or with a drop of water, or FNDs in aqueous solution in the optical trap, cf.
+Fig.
+1a)-c).
+The dry conditions facilitates the understanding of thermal phenomena, such as localized heating effects, and are intended as a clear reference to compare with conditions involving FNDs suspended in aqueous environments.
+We acknowledge that the dry condition itself lacks direct practical application in sensing contexts.
+For each set of experiments, we exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW, evaluated at the position of the sample.
+Fig.
+3 show five representative examples of variation in properties from individual FNDs as NIR power increases.
+As detailed below, data for FNDs on glass with a drop of water are not shown, but will be described.
+Figure 3: Representative examples of variation for individual nanodiamonds.
+(a) ODMR resonance frequency shift, relative to the expected resonance at 2.87 GHz, (b) fluorescence intensity ratio $\sigma_{0-}$, and (c) $T_{\rm relax}$ relaxometry times as functions of NIR laser power.
+Data for 5 random FNDs dry on glass (black stars) and trapped in water (gray circles) with 1064 nm NIR laser irradiation.
+Fig. 4 summarizes the results for 30 FNDs under dry conditions or in aqueous solution in the optical trap. Examples of the underlying measurements are provided in the Supplementary information.
+Figure 4: Summary statistics.
+a) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs trapped inside water.
+b) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs trapped inside water.
+c) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs dry on glass.
+d) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs dry on glass.
+Data for 30 FNDs exposed to 0 mW (green), 30 mW (orange), and 60 mW (red) of 1064 nm NIR laser irradiation.
+CW NIR for ODMR and spectra, pulsed NIR as in Fig 1d) for fluorescence relaxometry.
+In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and and no or much less pronounced trend for FNDs in the optical trap in water.
+To evaluate thermal effects, ODMR spectra were acquired in continuous-wave (CW) mode with constant intensities of both the 532 nm pump and 1064 nm trapping lasers and the microwave source, while sweeping the frequency of the microwave in steps of 0.3 MHz.
+Generally and for all FNDs, we observe a redshift in the resonance frequency with increasing CW NIR laser power.
+For FNDs dried on glass substrates, the resonance frequency shifted significantly to the lower frequencies.
+Taking 2.87 GHz as reference value for the zero field splitting, the resonance frequency shift increased in absolute value from $0.31\pm 0.54$ MHz at 0 mW to $5.83\pm 046$ MHz at 60 mW of CW NIR power.
+In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $0.84\pm 0.21$ MHz at 30 mW and $1.58\pm 0.33$ MHz at 60 mW.
+We also observe asymmetries in ODMR contrasts corresponding to spin transitions $\Delta m_{s}=\pm 1$, in correspondence with earlier results in literature for diamonds with ensembles of NV-centers Mittiga et al.
+(2018); Broadway et al.
+(2019); Bauch et al.
+(2020).
+Guided by these insights from literature, we interpret the asymmetries to be the result of local strain heterogeneity and uneven charge distributions near the diamond surface.
+To evaluate charge-state balance, we record the PL spectrum in the range between 550 nm to 745 nm with the spectrometer for each of the different values of the CW NIR laser power.
+Upon CW NIR laser light exposure, our PL measurements revealed a significant drop in the total PL intensity.
+The curves for FNDs on glass remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30 mW to 60 mW and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water.
+To quantify the changes in relative charge state emission intensities under CW NIR laser irradiation, we determine the ratio $\sigma_{0-}$ between NV0 and NV- emission intensities by integrating over the spectral ranges for NV0 (550–620 nm) and NV- (655–745 nm), respectively.
+Compared to the analysis used in Ref.
+Sow et al.
+(2020), we applied a normalization factor 19/14 to account for the difference in the relative spectral ranges:
+$\sigma_{0-}\equiv\left(\frac{\sum_{\lambda=550}^{620}I_{\lambda}}{\sum_{\lambda=655}^{745}I_{\lambda}}\right)\times\frac{19}{14}$ (1)
+For FNDs trapped with NIR powers of 30 mW and 60 mW, we observe an averaged similar emission ratios $\sigma_{0-}$ of $0.48\pm 0.09$.
+In contrast, for FNDs dried on glass, we measure an emission ratio of $0.43\pm 0.11$ without NIR irradiation that increases to $0.57\pm 0.15$ for a NIR laser power of 60 mW NIR.
+This is illustrated by the representative examples of fitted values for $T_{\rm relax}$ in Fig.
+3b) and the summary of the same for 30 individual FNDs in Fig.
+4 parts a) and c).
+Last, we conducted relaxometry experiments under different NIR laser power, to extract values for the corresponding $T_{\rm relax}$.
+For FNDs dried on glass substrates, we observed a significant decrease in the $T_{\rm relax}$ relaxation times with increasing NIR laser power.
+For FNDs trapped in water, the decrease in $T_{\rm relax}$ due to NIR irradiation was significantly less, and we hardly observe any visible change in the two relaxometry curves.
+Representative examples are shown in Fig.
+3c) with statistics of all results summarized in Fig.
+4, part b) and d).
+To study the significance of the observed effects, we performed an Analysis of Variance (ANOVA) test on the $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
+The ANOVA results are summarized in Table 1 (one-way Kruskal–Wallis test, $\alpha=0.05$).
+Table 1: ANOVA test results for the effect of NIR laser power on FNDs dried on glass substrates, on glass after adding water and trapped in water.
+For FNDs on glass, dry or with water, three groups were compared (NIR power of 0 mW, 30 mW or 60 mW) while for FNDs in the optical trap, only two groups (NIR power of 30 mW and 60 mW).
+The p-values indicate the significance of differences observed across NIR power levels for $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
+p-value
+FNDs dried on glass
+$T_{\rm relax}$ Relaxation Time <0.0001<0.0001
+ODMR Frequency Shift <0.0001<0.0001
+$\sigma_{0-}$ Ratio 0.0017
+FNDs on glass with water
+$T_{\rm relax}$ Relaxation Time $0.020$
+ODMR Frequency Shift $0.34$
+$\sigma_{0-}$ Ratio $0.019$
+FNDs Trapped in Water
+$T_{\rm relax}$ Relaxation Time 0.221
+ODMR Frequency Shift 0.073
+$\sigma_{0-}$ Ratio 0.371
+For FNDs dried on glass substrates, the very low p-values (p <0.01<0.01) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant.
+This indicates that NIR irradiation has a considerable impact on the NV centers’ properties in air.
+For FNDs on glass substrates after adding water, the p-values of $\simeq 0.02$ for $T_{\rm relax}$ and $\sigma_{0-}$ indicate statistical significance while the ODMR frequency shift, indicative of thermal effects, does not show statistically significant variation.
+In contrast, for FNDs trapped in water, the p-values are high (p >0.07>0.07), for all three quantities, suggesting that the differences observed across different NIR power levels are not statistically significant.
+Overall, as showed in Fig.
+3 and Fig.
+4, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts.
+Further, our results confirm that with specific NIR irradiation sequences and optimizing laser parameters, cf.
+Fig.
+1d) and Ref.
+Russell et al.
+(2018), potential detrimental effects of NIR on $T_{\rm relax}$ relaxation can be significantly mitigated.
+The rate equation simulation reveal that at low CW NIR laser powers, the population of NV- centers increases slightly, while the NV0 population decreases.
+This is attributed to enhanced recombination processes facilitated by the NIR photons, which convert NV0 centers back to NV-.
+Specifically, up to a CW NIR power of approximately 10 mW, the NV- population increases by about 8% compared to the case without CW NIR illumination.
+However, as the CW NIR power increases beyond this threshold, the trend reverses.
+The NV- population begins to decrease, and the NV0 population increases.
+At a CW NIR power of 60 mW, the NV- population decreases by approximately 10% from its maximum value, while the NV0 population increases correspondingly.
+This inversion is due to the saturation of recombination processes and the dominance of ionization processes at higher CW NIR powers, which ionize NV- centers back to NV0.
+Consequently, the total fluorescence intensity from the NV centers first increases then decreases with increasing CW NIR power (Fig.
+5).
+This decrease is primarily due to the reduced population of NV- centers and the lower fluorescence quantum yield of NV0 centers.
+The fluorescence from NV- centers dominates the emission under low CW NIR power, but as the CW NIR power increases, the contribution from NV0 centers becomes more significant.
+At low CW NIR powers, the $\sigma_{0-}$ ratio is low due to the predominance of NV- fluorescence.
+As the CW NIR power increases, the ratio rises slowly, indicating a relative increase in NV0 emission.
+This change in the fluorescence ratio affects the overall spectral characteristics and can influence the detection sensitivity in applications relying on NV- fluorescence.
+The trends observed in the simulation align with the experimental results.
+The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity (Fig.
+5) and $\sigma_{0-}$ ratio (Fig.
+4a,c)) observed experimentally.
+Figure 5: Total fluorescence intensity from NV centers as a function of NIR laser power.
+Black: Theory, Red: Experimental observations with errorbars, obtained as a statistical average over results from 10 different FNDs with $\sigma_{0-}$ within the range of $\sigma_{0-}=0.42\pm 0.04$.
 […]
-In Water: The temperature increase of the FND is minimal due to the high thermal conductivity and heat capacity of water. The simulation indicates that during the 500 ns NIR pulse:
-  • The maximum temperature rise ($\Delta T_{\text{max}}$) is approximately 2 K above the ambient temperature.
-  • The temperature increases steadily during the laser pulse and rapidly returns to ambient temperature within nanoseconds after the laser is turned off.
-In Air: The temperature increase is more pronounced because air has lower thermal conductivity and heat capacity. The simulation results show:
-  • The maximum temperature rise reaches approximately 41 K above the ambient temperature during the 500 ns NIR pulse.
-  • The temperature increases during the laser pulse and decreases more slowly than in water after the laser is turned off.
-The simulation results align with experimental observations where pulsed NIR illumination causes negligible thermal effects in aqueous environments, cf.
+The results of the simulation of thermal effects align with experimental observations where NIR illumination causes notable thermal effects in air and much less in aqueous environments, cf.
 Fig.
-7.
-In experiments conducted in air, temperature increases contribute to shifts in ODMR frequency and should be accounted for.
-These shifts correspond to a 2 MHz shift in the ODMR resonance frequency which is what we experimentally observe.
+3-Fig.
+4.
 A direct comparison of Fig.
-6c)–d) with Fig.
-6a)–b) shows that the environment (air versus water) is the controlling parameter.
+4c)–d) with Fig.
+4a)–b) shows that the environment (air versus water) is the controlling parameter when it comes to the thermal response.
 In air, the ODMR resonance shifts by $5.8\pm 0.5$ MHz and $T_{\rm relax}$ falls by $(44\pm 6)\%$ when the CW NIR power is raised from 0 to 60 mW, whereas in water the corresponding changes are only $1.6\pm 0.3$ MHz and $(2\pm 5)\%$, respectively.
 By contrast, the NV0/NV- fluorescence intensity ratio $\sigma_{0-}$ responds similarly to NIR power in both media (Fig.
-6a,c; projection on the ordinate axis), confirming that charge-state kinetics are governed by photon flux rather than heat removal.
-For completeness we note that NV- centres possess a weak intersystem transition between the metastable singlet states $^{1}E\,(p_{6})\!\rightarrow\!{}^{1}A_{1}\,(p_{5})$ with a zero-phonon line at $\lambda_{\mathrm{ZPL}}=1042$ nm and a room-temperature homogeneous width of only a few GHz ($\lesssim 10^{-4}$ nm) [Acosta2010, Rogers2015].
+4a,c; projection on the ordinate axis), confirming that charge-state kinetics are governed by photon flux rather than heat exchange.
+For completeness we note that NV- centres possess a weak intersystem transition between the metastable singlet states $^{1}E\,(p_{6})\!\rightarrow\!{}^{1}A_{1}\,(p_{5})$ with a zero-phonon line at $\lambda_{\mathrm{ZPL}}=1042$ nm and a room-temperature homogeneous width of only a few GHz ($\sim 10^{-4}$ nm) Rogers et al.
+(2015).
 Our trapping beam at 1064 nm is therefore red-detuned by $\Delta\lambda=22$ nm, corresponding to $\Delta\nu\simeq 6$ THz ($\simeq$  25 meV), i.e.
 $\sim 10^{3}$ times the natural linewidth and well outside the vibrational sideband of the singlet manifold.
-Values for the measured absorption cross-sections decrease from $\sigma\!\approx\!2\times 10^{-19}$ cm2 on resonance to below $10^{-21}$ cm2 at this detuning [Acosta2010], so direct optical pumping of the singlet transition by the 1064 nm light is negligible.
-The effects reported in this work (PL quenching, ODMR shifts, $T_{\rm relax}$ variations) therefore arise from photothermal heating and photo-ionisation dynamics, not from resonant excitation of the 1042 nm line.
-IV Sensing Capabilities of Trapped FNDs
-As a final step in the comprehensive investigations, this section explores the potential of optically trapped FNDs with NV centers for sensing applications.
-Experiments focused on three key parameters: Gd3+ concentration, pH, and temperature.
-Measurements included ODMR, fluorescence relaxation, and optical spectrum analysis under variation of the three parameters mentioned above.
+The effects reported in this work (PL quenching, ODMR shifts, $T_{\rm relax}$ variations) therefore most likely arise from photothermal heating and photo-ionisation dynamics, not from resonant excitation of the 1042 nm line.
+3.1 Sensing Capabilities of Trapped FNDs
+In addition to the multi-modal characterization presented above, the potential of optically trapped FNDs with NV centers for sensing applications were addressed.
+Experiments focused on three key parameters: Gd3+ concentration, pH, and temperature, and the same three measurement modalities (PL spectroscopy, ODMR, and optical relaxometry) as before.
 Results are summarized in Fig.
-10.
+6.
 Figure 6: Variation in three sensing parameters when subject to controlled variation of paramagnetic species, pH, and temperature.
@@
 The red squares emphasize the optimal measurement method for each sensing parameter.
-IV.1 pH Sensing
+The temperature response of nanodiamonds with NV centers was evaluated under three conditions: FNDs on glass substrates with no NIR laser irradiation, trapped FNDs with no temperature increase, and trapped FNDs with temperature increases of +1 K and +2 K.
+As expected, a temperature-dependent shift in the ODMR resonance frequency was observed, with lower frequencies at higher temperatures (Fig.
+6c)).
+This shift is attributed to the well-known temperature dependence of the NV center’s zero-field splitting parameter $D$, which decreases at a rate of approximately $74~\text{kHz/K}$ Acosta et al.
+(2010).
+On the contrary, no significant changes were detected in $T_{\rm relax}$ values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig.
+6i)).
+Further, no significant changes in $\sigma_{0-}$ were observed with temperature, suggesting that charge state dynamics are minimally affected by temperature within the range investigated here (Fig.
+6f)).
+ODMR’s temperature sensitivity highlights its utility for high-resolution nanoscale thermometry.
+The absence of significant $T_{\rm relax}$ or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements.
 The impact of environmental pH on the properties of NV centers was evaluated using buffer solutions with pH values of 1, 7, and 13.
@@
 Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
-  • Fluorescence relaxation: A decrease in $T_{\rm relax}$ values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig.
-10.h).
-This is consistent with reports that high pH conditions stabilize spin relaxation dynamics [Fujisaku2019, Rendler2017].
-  • Optical Spectrum: A decrease in the $\sigma_{0-}$ ratio with increasing pH (Fig.
-10.e and Fig.
-11) suggests that alkaline conditions favor the NV- charge state.
+No significant changes in ODMR resonance frequency (Fig.
+6b)) were detected across the tested pH range, indicating minimal impact of pH on electronic spin states.
+A decrease in $T_{\rm relax}$ values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig.
+6h)).
+This is consistent with reports that high pH conditions stabilize spin relaxation dynamics.
+A decrease in the $\sigma_{0-}$ ratio with increasing pH (Fig.
+6e)) and Fig.
+7) suggests that alkaline conditions favor the NV- charge state.
 It may be attributed to the deprotonation of surface functional groups such as carboxyl (–COOH) and hydroxyl (–OH) groups present on the nanodiamond surface.
-Deprotonation of these groups under alkaline conditions generates negatively charged surface sites, thereby enhancing electron donation to nearby NV centers and favoring the stabilization of the NV- charge state [karaveli2016].
+Deprotonation of these groups under alkaline conditions generates negatively charged surface sites, thereby enhancing electron donation to nearby NV centers and favoring the stabilization of the NV- charge state Karaveli et al.
+(2016).
 The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (Fig.
-11) shows a clear advantage of using FNDs in suspension for sensing.
-  • ODMR: No significant changes in resonance frequency (Fig. 10.b) were detected across the tested pH range, indicating minimal impact of pH on electronic spin states.
-The $\sigma_{0-}$ ratio’s dependence on pH underscores the importance of surface chemistry in modulating charge state dynamics, especially with FNDs in suspension.
+7) shows a clear advantage of using FNDs in suspension for sensing.
+Another advantage of the optical trap is that effects of individual variation between FNDs from a batch is reduced since measurements can be made relative to an initial measurement with the same FND.
 To further explore the dependence of the NV charge state on pH, Fig.
@@
 This enhanced responsiveness highlights the improved surface accessibility and chemical exchange in solution, reinforcing the utility of optically trapped FNDs for pH nanosensing.
-The pronounced increase in NV center fraction under alkaline conditions (lower $\sigma_{0-}$) confirms the surface deprotonation hypothesis, while acidic conditions suppress this stabilization.
-Overall, these results validate that $\sigma_{0-}$ serves as a reliable and sensitive optical proxy for probing local pH variations in biological or environmental settings.
-IV.2 Paramagnetic Species Sensing
+Overall, the results confirm that $\sigma_{0-}$ serves as a reliable and sensitive optical proxy for probing local pH variations in biological or environmental settings.
 The sensitivity of NV centers to paramagnetic gadolinium ions (Gd3+) was investigated through measurements at 3 different values of the concentration.
@@
 Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
-  • Fluorescence relaxation: A significant decrease in $T_{\rm relax}$ values was observed with increasing Gd3+ concentration (Fig.
-10.g).
-This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment [Martinez2020, Rendler2017].
-  • Optical Spectrum: A decrease in $\sigma_{0-}$ was observed (Fig.
-10.d and Fig.
-12).
-This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 [Rondin2010].
+No significant shifts or changes in ODMR resonance frequency were detected, suggesting limited interaction between Gd3+ ions and the electronic spin states of the NV centers under the tested conditions (Fig.
+6a)).
+A decrease in $\sigma_{0-}$ was observed (Fig.
+6d) and Fig.
+8).
+This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 Rondin et al.
+(2010).
 The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (fig.
-12) shows a clear advantage of using FNDs in suspension for sensing.
-  • ODMR: No significant shifts or changes in resonance frequency were detected, suggesting limited interaction between Gd3+ ions and the electronic spin states of the NV centers under the tested conditions (Fig.
-10.a).
-These findings demonstrate the potential of NV centers for detecting paramagnetic species like Gd3+.
-The sensitivity saturates at higher concentrations due to the finite interaction volume and saturation of surface sites for Gd3+ binding.
+8) shows a clear advantage of using FNDs in suspension for sensing.
+Finally, a significant decrease in $T_{\rm relax}$ values was observed with increasing Gd3+ concentration (Fig.
+6g)).
+This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment Tetienne et al.
+(2013); Rendler et al.
+(2017); Martinez et al.
+(2020).
+These findings demonstrate the potential of optically trapped nanodiamonds with NV centers for detecting paramagnetic species like Gd3+.
 The fluorescence relaxometry provides a robust method for quantifying Gd3+ in biological and environmental contexts.
@@
 For completeness, we observe that by adding Gd3+ to a concentration of 0.1 M, the pH value in the solution increases slightly (by of order 1-2 pH units) which is consistent with our previous observations on the effect of pH on the fluorescence of FNDs.
-IV.3 Temperature Sensing
-The temperature response of NV centers was evaluated under three conditions: FNDs on glass substrates with no NIR laser irradiation, trapped FNDs with no temperature increase, and trapped FNDs with temperature increases of +1 K and +2 K.
-  • ODMR: A temperature-dependent shift in the resonance frequency was observed, with lower frequencies at higher temperatures (Fig.
-10.c).
-This shift is attributed to the well-known temperature dependence of the NV center’s zero-field splitting parameter $D$, which decreases at a rate of approximately $74~\text{kHz/K}$ [Acosta2010].
-  • Fluorescence relaxation: No significant changes were detected in $T_{\rm relax}$ values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig.
-10.i).
-  • Optical Spectrum: No significant changes in $\sigma_{0-}$ were observed with temperature, suggesting that charge state dynamics are minimally affected by thermal fluctuations (Fig. 10.f).
-ODMR’s temperature sensitivity highlights its utility for high-resolution nanoscale thermometry.
-The absence of significant $T_{\rm relax}$ or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements
-V Discussion
+4 Discussion
 The shift in the ODMR resonance frequency with increasing NIR power, as shown in Fig.
-4, Fig.
-6, and Fig.
-7a), suggests an increase in temperature within the FNDs.
-This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately 74-74 kHz/K [Acosta2010, Toyli2012].
+3a), and Fig.
+4, suggests an increase in temperature within the FNDs.
+This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately 74-74 kHz/K Acosta et al.
+(2010); Toyli et al.
+(2012) and corresponds with previous suggestions of the heating effect of NIR light Lai et al.
+(2013).
 However, the extent of the thermal effect is significantly influenced by the surrounding medium’s properties.
 Media with higher thermal conductivity and heat capacity, such as water, dissipate heat more efficiently than air, resulting in smaller temperature increases in the FNDs.
-This is supported by our temperature simulations, which show minimal temperature rise in water compared to air during NIR illumination.
-Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts with increasing NIR power, as shown in Fig.
-6 and Fig.
-7.
+This is supported by our temperature simulations, which show much lower temperature rise in water compared to air during NIR illumination.
+Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts with increasing NIR power.
 In addition to thermal effects, the NIR laser alters the charge-state dynamics of the NV centers.
@@
 This inversion is attributed to the saturation of recombination processes and increased ionization rates at higher NIR powers.
-The total fluorescence intensity decreases with increasing NIR power, as shown in Fig.
-9.
-This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers [Chapman2010].
+As a result, the total fluorescence intensity decreases with increasing NIR power, as shown in Fig.
+5.
+This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers Chapman and Plakhotnik (2011).
 The increase in $\sigma_{0-}$ with higher NIR power, depicted in Fig.
-6 and Fig.
-7, further confirms the shift in charge state distribution.
+3b) and Fig.
+4, further confirms the shift in charge state distribution.
 Additionally, the increased NV0 background fluorescence dilutes the ODMR contrast since NV0 centers do not exhibit spin-dependent fluorescence variations.
-This is reflected in the decreased ODMR contrast observed in our measurements (Fig.
-4).
-Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{\rm relax}$ times [Aslam2013, Goudarzi2019].
+This is reflected in the decreased ODMR contrast observed in our measurements (data not shown).
+Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{\rm relax}$ times Aslam et al.
+(2013).
+With an all-optical protocol for relaxometry as used here, the results for $T_{\rm relax}$ presented inherently involves potential inaccuracies due to photoionization induced by the green laser itself, thus affecting the NV- charge states.
+To address this limitation, measurement protocols incorporating microwave pulses explicitly designed to disentangle spin dynamics from photoionization effects have been developed.
+Examples include the microwave $\pi$-pulse method introduced by Ref.
+Jarmola et al.
+(2012), as well as alternative sequences described in Ref.
+Cardoso Barbosa et al.
+(2023).
 Our results indicate that although both thermal effects and charge-state dynamics significantly impact the properties of the NV centers under NIR illumination, their relative contributions depend on the environmental conditions.
@@
 The increase in the $\sigma_{0-}$ ratio and the decrease of fluorescence relaxometry times with NIR power without significant ODMR frequency shifts (Fig.
-6) indicate that charge-state transitions are the primary mechanism affecting the fluorescence and spin properties in aqueous environments.
+4) indicate that charge-state transitions are the primary mechanism affecting the fluorescence and spin properties in aqueous environments.
 While thermal effects and charge-state dynamics provide a comprehensive explanation for our observations, alternative mechanisms should be considered.
 Multi-photon absorption processes could directly excite the NV centers without significant heating, yet, the relationship between ODMR frequency shift and NIR power (Fig.
-6) is contrary to this explanation.
+4) is contrary to this explanation.
 Moreover, the significant differences in the effects of NIR irradiation between FNDs in air and water emphasize the critical role of the surrounding medium.
 The statistical analysis using ANOVA (Table 1) confirms that the differences observed in FNDs dried on glass are statistically significant, while those in FNDs trapped in water are not.
+It should also be emphasized that while our results are consistent with an interpretation of ODMR frequency shift being proportional to a temperature shift, the interpretation may be more complex since changes in local electrical field also implies shifts of the ODMR frequency Dolde et al.
+(2011), an effect that we cannot completely exclude in our experiments.
+Similarly, literature suggests NIR-induced dark state transitions with strong NIR pulses Geiselmann et al.
+(2013); Cao et al.
+(2025b, a), and including these extra states in our theoretical model could provide more detailed understanding of our results.
 In our opinion, the data presented in this work establish a clear hierarchy of influences: (i) the thermal conductivity and heat capacity of the surrounding medium dictate the magnitude of the NIR-induced temperature rise and therefore the thermally induced shifts in ODMR frequency and the shortening of $T_{\rm relax}$; (ii) the NIR photon flux sets the rate of photo-ionisation and recombination, which in turn controls the NV0/NV- balance and the fluorescence contrast.
-Optical trapping itself plays a mainly mechanical role-holding a single FND at a fixed position—while potential effects on spin and charge-state dynamics depend on the wavelength of the trapping laser.
+Optical trapping itself plays a mainly mechanical role–holding a single FND at a fixed position—while potential effects on spin and charge-state dynamics depend on the wavelength of the trapping laser.
 This distinction is important for two reasons.
@@
 Second, it generalises our findings to intracellular sensing, where the cytoplasm’s thermal properties resemble those of water: the modest $\lvert\Delta f\rvert<2$ MHz and unchanged $T_{\rm relax}$ observed in vitro set an upper bound on in-cell NIR heating, validating the use of trapped FNDs for quantitative biosensing.
-The results presented in Section IV confirm that optical trapping of FNDs provide a valuable add-on that allows to enhance the application potential of biosensing with NV-centers in nanodiamonds.
+The results presented in the sensing section confirm that optical trapping of FNDs provide a valuable add-on that allows to enhance the application potential of biosensing with NV-centers in nanodiamonds.
 In particular, Fig.
-10a-c) demonstrates that the frequency shift in ODMR is a good stand-alone measure for a temperature shift also for optically trapped FNDs, while parts f) and i) illustrate that changes in the surrounding temperature within a range that is relevant for biological systems does not appear to alter the readings provided by the fluorescence intensity ratio $\sigma_{0-}$ or the relaxation time, $T_{\rm relax}$.
-VI Conclusion
-This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive understanding of the interplay between thermal effects, charge-state dynamics, and their combined influence on sensing capabilities.
+6a-c) demonstrates that the frequency shift in ODMR is a good stand-alone measure for a temperature shift also for optically trapped FNDs, while parts f) and i) illustrate that changes in the surrounding temperature within a range that is relevant for biological systems does not appear to alter the readings provided by the fluorescence intensity ratio $\sigma_{0-}$ or the relaxation time, $T_{\rm relax}$.
+5 Conclusion
+This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive overview of the interplay between thermal effects, charge-state dynamics, and their combined influence on sensing capabilities.
 By employing an integrated experimental approach combining photoluminescence spectroscopy, optically detected magnetic resonance (ODMR), and fluorescence relaxometry, this study elucidates how varying NIR laser powers impact the fluorescence, spin properties, and stability of NV centers in diverse environmental conditions.
@@
 Long-term stability under extended NIR irradiation remains an open question, particularly regarding potential degradation of fluorescence and spin properties.
+The role of surface functionalization in enhancing NV center stability and sensing accuracy requires further exploration, as does the potential for alternative NIR wavelengths or advanced laser modulation techniques to improve performance.
 The effects of complex biological environments, including interactions with proteins, ions, and other cellular components, are yet to be fully understood.
-The role of surface functionalization in enhancing NV center stability and sensing accuracy requires further exploration, as does the potential for alternative NIR wavelengths or advanced laser modulation techniques to improve performance.
 Theoretical modeling, while providing valuable insights into charge state dynamics and thermal effects, must be expanded to include the complexities of real-world applications.
 Future work could explore the integration of complementary sensing techniques, such as fluorescence lifetime imaging or advanced magnetic resonance methods, to provide a a real-time sensing protocol inside cells by using optically trap FNDs in the cellular medium.
-Furthermore, efforts to miniaturize and automate the experimental setup, enabling in vivo applications, would be a critical step in translating this research into practical diagnostic tools.
 Acknowledgments
@@
 We further acknowledge insights provided through the student project work of Kristian Lambertsen and Nicolai Kongstad, as well as Suyash Amzare, including an original sketch of the experimental setup providing the basis for Fig.
-2.
+9.
 Appendix
-We specify in Table 2 all part numbers in the optical setup, Fig. 2a).
-Table 2: Optical setup component
+Figure 9: Detailed illustration of the optical setup with the most important lasers, optical lenses, filters and equipment. The components are specified in Table 2.
+We specify in Table 2 all part numbers in the optical setup, Fig. 9.
+Table 2: Components of the optical setup, Fig. 9. Filters at position 23 and 26 are switchable.
 Ref. Number Component Model
@@
 7 Iris Aperture ThorLabs ID12Z/M
-8 IR disk ThorLabs VRC4CPT
-9 Beam expander
-10 Right-angle prism ThorLabs DFM1/M
+9 Beam expander Part of Thorlabs OTKB/M
+10 Dicroic mirror DMSP805R (in ThorLabs DFM1/M)
 11 Green laser Verdi G-Series - Coherent
@@
 20 Lens ThorLabs LA1509 - A
-21 Right-angle prism ThorLabs DFM1/M
+21 Dicroic mirror Semrock NFD01-532-25x36 (in Thorlabs DFM1/M)
 22 Lens ThorLabs LA1484 - A
-23 Filters Bightline Basic 460/60 - ThorLabs MF630-69 Emission filter
+23 Filters Brightline Basic 460/60 - ThorLabs MF630-69 Emission filter
 24 Camera Teledyne Prime 95B
 25 Filter ThorLabs NE20A
-26 Filters Bightline Basic FF01-630/52 - ThorLabs FELH 0550
+26 Filters Brightline Basic FF01-630/52 - ThorLabs FELH 0650
 27 Objective lens Olympus PLN 10X
@@
 34 Quadrant photoelectric detector ThorLabs PDQ80A
-35 White lamp
+35 White lamp Part of Thorlabs OTKB/M
 36 Spectrometer Ocean Optics QE Pro
@@
 We applied the relevant values from Ref.
-[Qian2022] and combined them with parameters from [Meirzada2018] for the four ionization and recombination rates as these authors use the same laser wavelengths as us in their protocol.
-Below, in Table 3, we provide the values used in the simulations of Section III.4.
+Qian et al.
+(2022) and combined them with parameters from Meirzada et al.
+(2018) for the four ionization and recombination rates as these authors use the same laser wavelengths as us in their protocol.
+Below, in Table 3, we provide the values used in the simulations of Section 2.5.
 Table 3: Parameters used in the simulation
@@
 $K_{rIR}$ 3.17 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Recombination rate (IR laser)
-laser_green_power 0.7 mW Green laser power
+$I_{G}$ 0.7 mW Green laser power